### v1

# Refining Concentration for Gaussian Quadratic Chaos

Abstract
We visit and slightly modify the proof of Hanson-Wright inequality (HW inequality) for concentration of Gaussian quadratic chaos where we are able to tighten the bound by increasing the absolute constant in its formulation from its largest currently known value of 0.125 to at least 0.145 in the symmetric case. We also present a sharper version of the so-called Laurent-Massart inequality (LM inequality) through which we are able to increase the absolute constant in HW inequality from its largest currently available value of $0.134$ due to LM inequality itself to at least $0.152$ in the positive-semidefinite case. Generalizing HW inequality in the symmetric case, we derive a sequence of concentration bounds for Gaussian quadratic chaos indexed over $m=1,2,3,\cdots$ that involves the Schatten norms of the underlying matrix. The case $m=1$ reduces to HW inequality. These bounds exhibit a phase transition in behaviour in the sense that $m=1$ results in the tightest bound if the deviation is smaller than a critical threshold and the bounds keep getting tighter as the index $m$ increases when the deviation is larger than the aforementioned threshold. Finally, we derive a concentration bound that is asymptotically tighter than HW inequality both in the small and large deviation regimes.
I Introduction
I-A Summary of prior art
This paper concerns the Gaussian version of Hanson-Wright (HW) inequality [1], a concentration of measure result that presents an upper bound on the tail probability for a quadratic form in a random vector of independent sub-Gaussian random variables also known as order 2 chaos. For fixed integer $n\geq 1$, let $\underline{\boldsymbol{x}}$ be a vector of $n$ independent standard Gaussian random variables and $A$ be an arbitrary $n\times n$ matrix. Denote the difference between $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}$ and its expected value $\mathbb{E}[\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}]=\mathrm{Tr}(A)$ by $\boldsymbol{\Delta}$, i.e.,
$\boldsymbol{\Delta}=\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}-\mathrm{Tr}(A).$ (1)
HW inequality states that for every $t>0$, [^1] Replacing $A$ by $-A$ in (2) results in the same upper bound as in (2) on the left tail probability $\Pr(\boldsymbol{\Delta}<-t)$.
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\Big{(}-\kappa\min\Big{\{}\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\Big{\}}\Big{)},$ (2)
where $\|A\|_{2}$ is the Hilbert-Schmidt norm of $A$, $\|A\|$ is the operator norm of $A$ and $\kappa$ is an absolute constant that does not depend on $n$, $A$ and $t$. Without loss of generality, we assume $A$ is symmetric throughout the paper. This can be easily checked by noting that $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}=\underline{\boldsymbol{x}}^{T}\tilde{A}\underline{\boldsymbol{x}}$ where $\tilde{A}=\frac{1}{2}(A+A^{T})$ is symmetric and $\|\tilde{A}\|_{2}\leq\|A\|_{2}$ and $\|\tilde{A}\|\leq\|A\|$. These inequalities follow by applying the triangle inequality for norms and noting that $\|A\|_{2}=\|A^{T}\|_{2}$ and $\|A\|=\|A^{T}\|$. In the original paper [1], $\|A\|$ was defined as the operator norm of the matrix whose entries are absolute values of entries of $A$. Recently, reference [2] shows that the inequality indeed holds as stated in (2) with $\|A\|$ being the operator norm of $A$ itself. An explicit value for $\kappa$ is missing or at least hard to identify in [1, 2]. More recently, the textbook[3] gives the explicit value $\kappa=\frac{1}{8}=0.125$ at the end of the proof for Theorem B.8 on page 304. This value for $\kappa$ can also be inferred form inequality (3.28) on page 120 in [4] given by
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\Big{(}-\frac{t^{2}}{4(\|A\|_{2}^{2}+\|A\|t)}\Big{)}.$ (3)
It is stated in [4] that (3) holds for arbitrary symmetric matrix $A$. However, there is a mistake in the proof. Inequality $\frac{1}{2}(-\ln(1-2x)-2x)\leq\frac{x^{2}}{1-2x}$ in the middle of page 120 does not hold for $x<0$. Nonetheless, inequality (3) certainly holds when $A$ is positive-semidefinite. Loosening the bound by writing $\|A\|_{2}^{2}+\|A\|t\leq 2\max\{\|A\|_{2}^{2},\|A\|t\}$, one arrives at HW inequality with $\kappa=0.125$. For a positive-semidefinite matrix $A$, a key result in the literature on concentration of Gaussian quadratic chaos is inequality (4.1) on page 1325 in [5] given by
$\Pr(\boldsymbol{\Delta}>2\|A\|_{2}\sqrt{t}+2\|A\|t)\leq\exp(-t),$ (4)
for all $t>0$. We will refer to (4) as Laurent-Massart inequality (LM inequality). We will show in Section  that LM inequality implies HW inequality at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
I-B Contributions and main results
The contributions in this paper are threefold:
  1. 1.
    We visit and slightly modify the proof given in [3] for HW inequality. Reference [3] uses the bound
    $-\ln(1-x)\leq x+x^{2},\,\,\,|x|\leq\frac{1}{2},$ (5)
    which results in $\kappa=0.125$. Instead, we consider the bound
    $-\ln(1-x)\leq x+ax^{2},\,\,\,|x|\leq b,$ (6)
    for $a,b>0$. For given $a>\frac{1}{2}$, we determine the largest $b$ such that (6) holds. As a result, we are able to increase $\kappa$ from $0.125$ to at least $0.145$ for arbitrary symmetric matrix $A$. More precisely, we have the following proposition.
    Proposition 1.
    For $10<b<$ 1 and integer $m\geq 1$ define
    $\theta_{m}(b)=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}.$ (7)
    If $A$ is symmetric, then HW inequality holds with
    $\kappa=\frac{b^{*}}{4}\approx 0.1457,$ (8)
    where $b^{*}$ is the unique solution to the equation $12b\theta_{1}(b)=$ 1.
    Proof.
    See Section III.∎
  2. 2.
    In the positive-semidefinite case, we present a sharper version of LM inequality. To prove (4), reference [5] relies on the bound
[…]
    Instead, we consider
[…]
    for $a,b>0$. We show that for every $\frac{2}{3}<a\leq 1$, this inequality holds for $b$ at least as large as $\frac{3a-2}{a(2a-1)}$. As a result, we derive an improved LM inequality as presented in the next proposition.
    Proposition 2.
    Let $A$ be a positive-semidefinte matrix and
    $\alpha=\|A\|,\,\,\,\,\,\beta=\|A\|_{2}^{2}.$ (11)
    For $\frac{2}{3}<a\leq 1$, let
    $b=\frac{3a-2}{a(2a-1)},\,\,\,\,\,c=\frac{3a-2}{2(1-a)^{2}}.$ (12)
    Then
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda(t,a)},$ (13)
    where
    $\Lambda(t,a)=\left\{\begin{array}[]{cc}-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\Big{(}\sqrt{1+\frac{2a\alpha t}{\beta}}-1\Big{)}&\frac{\alpha t}{\beta}\leq c\\ -\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}&\frac{\alpha t}{\beta}>c\end{array}\right..$ (16)
    Proof.
    See Section IV. ∎
    Choosing $a=1$ in (13) results in LM inequality. In Section  we will minimize the upper bound on the right side of (13) in terms of $a$ and compare the resulting inequality with LM inequality to see the improvement. As we will exhibit in Section II, the “optimum” $a$ is the unique root inside the interval $(\frac{2}{3},1)$ of a quintic polynomial equation which must be solved numerically. We are able to approximate this unique root by an analytic formula where the approximation error is shown to be less that $0.005$ regardless of the matrix $A$ and the deviation $t>0$. For our last result in this part, we establish the next proposition which is a consequence of Proposition 2.
    Proposition 3.
    Let $A$ be positive-semidefinite. The inequality in (13) implies HW inequality at best with
    $\kappa=\frac{9-\sqrt{17}}{32}\approx 0.1524,$ (17)
    which is achieved for
    $a=\frac{7-\sqrt{17}}{4}.$ (18)
    Moreover, LM inequality in (4) implies HW inequality with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
    Proof.
    See Section V. ∎
  3. 3.
    We explore beyond HW inequality for arbitrary symmetric matrix $A$ by considering an extension of (6), i.e., an inequality of the form
    $-\ln(1-x)\leq x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+a|x|^{m+1},\,\,\,\,\,|x|\leq b,$ (19)
    for arbitrary integer $m\geq 1$ and $a,b>0$. We show that for every integer $m\geq 1$ and $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (19) holds is $\theta_{m}(b)$ as defined in (7). As a consequence, we obtain a sequence of concentration bounds for Gaussian quadratic chaos which involve the Schatten norms [^2] See Subsection I.C for definition of the Schatten $p$-norm. of $A$.
    Proposition 4.
    Let $10<b<$ 1, $m\geq 1$ be an integer and $\|A\|_{m+1}$ be the Schatten norm of a symmetric matrix $A$ of order $m+1$. Then
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{\Omega_{m}(t,b)},$ (20)
    where
    $\Omega_{m}(t,b)=\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k}-\kappa_{m}(b)\min\bigg{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg{\}}$
    and
    $\kappa_{m}(b)=\frac{1}{2}\Big{(}1-\frac{1}{m+1}\Big{)}\min\Big{\{}\frac{1}{\big{(}(m+1)\theta_{m}(b)\big{)}^{\frac{1}{m}}},b\Big{\}}$ (21)
    with $\theta_{m}(b)$ as defined in (7).
    Proof.
    See Section VI. ∎
    Choosing $m=1$ in (20) recovers HW inequality in (2) with $\kappa=\kappa_{1}(b)$. The Schatten norm $\|A\|_{m+1}$ should not be mistaken with the $L_{p}$-norm of $A$ with $p=m+1$ unless $m=1$ in which case $\|A\|_{2}$ is the Hilbert-Schmidt norm of $A$. The upper bound $e^{\Omega_{m}(t,b)}$ in (20) is to be minimized over $10<b<$ 1. In general, the sequence of upper bounds $\inf_{b\in(0,1)}e^{\Omega_{m}(t,b)}$ for $m=1,2,3,\cdots$ is not monotone in $m$ for a given matrix $A$ and $t>0$. However, it undergoes a phase transition in the sense that if $t$ is sufficiently small, then it is the tightest for $m=1$ and if $t$ is sufficiently large, then it keeps getting tighter as $m$ increases. In Section II, we study these bounds more closely both through numerical examples and analytically. In particular, two corollaries of Proposition 4 are stated. One offers a looser version of the upper bound in (20) which is in a sense more convenient to use for large enough values of $t$ as it will not involve minimization over $b$ anymore. The other looks into the limiting value of (20) as $m$ grows large. Finally, we are able to present a twin inequality to HW inequality which is tighter than HW inequality for sufficiently small $t$ and sufficiently large $t$. This bound depends on all of $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$.
    Proposition 5.
    Let $A$ be a symmetric matrix and
    $\alpha=\|A\|,\,\,\,\beta=\|A\|_{2}^{2},\,\,\,\gamma=\|A\|_{3}^{3}.$ (22)
    Also, let
    $\eta_{1}(t)=\frac{1}{12\gamma^{2}}(\sqrt{\beta^{2}+4\gamma t}-\beta)(\beta^{2}+8\gamma t-\beta\sqrt{\beta^{2}+4\gamma t}).$ (23)
    and
    $\eta_{2}(t)=\frac{t}{\alpha}-\frac{3\beta}{4\alpha}\min\Big{\{}\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t-\beta}}{2\gamma}\Big{\}}.$ (24)
    Then
    $\Pr(\boldsymbol{\Delta}>t)\leq\exp\big{(}-\kappa^{\prime}\min\{\eta_{1}(t),\eta_{2}(t)\}\big{)},$ (25)
    where $\kappa^{\prime}=\frac{b^{*}}{3}\approx 0.187$ with $b^{*}$ as the unique solution in $b$ to the equation $\frac{4}{\sqrt{3}}b\sqrt{\theta_{2}(b)}=1$ and $\theta_{2}(b)$ is defined in (7).
    Proof.
    See section VII. ∎
    The bound in (25) is dubbed a twin to HW inequality in the sense that $\eta_{1}(t)=\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})$ as $t\to 0$ and $\eta_{2}(t)=\frac{t}{\|A\|}+O(1)$ as $t\to\infty$. Note that $\kappa^{\prime}\approx 0.187$ is slightly larger than the constant $\kappa\approx 0.145$ in HW inequality.
I-C Notations
The (real) eigenvalues of an $n\times n$ symmetric matrix $A$ are denoted by $\lambda_{1},\cdots,\lambda_{n}$. The operator norm of $A$ is
$\|A\|=\max_{1\leq i\leq n}|\lambda_{i}|,$ (26)
The Schatten $p$-norm of $A$ for $p\geq 1$ is defined by
$\|A\|_{p}=\big{(}\sum_{i=1}^{n}|\lambda_{i}|^{p}\big{)}^{1/p}$ (27)
The case $p=2$ gives the Hilbert-Schmidt norm of $A$. If $A$ is positive-semidefinite, then
$\|A\|_{p}=(\mathrm{Tr}(A^{p}))^{1/p},$ (28)
where $\mathrm{Tr}(\cdot)$ is the trace operator. Random variables are denoted by boldfaced letters such as $\boldsymbol{x}$ with expectation $\mathbb{E}[\boldsymbol{x}]$.
II Analysis of the bounds in Subsection I.B
III Proof of Proposition 1
The indispensable starting point in proving HW inequality is the exponential Markov inequality which gives
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-st}\mathbb{E}[e^{s\boldsymbol{\Delta}}],$ (29)
where $s\geq 0$ is arbitrary. Let $A=UDU^{T}$ be the eigenvalue decomposition of $A$ where $U$ is an orthogonal matrix and the diagonal matrix $D$ carries the (real) eigenvalues $\lambda_{1},\cdots,\lambda_{n}$ of $A$. Then
$\boldsymbol{\Delta}=\sum_{i=1}^{n}\lambda_{i}(\boldsymbol{\xi}^{2}_{i}-1),$ (30)
where $\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{n}$ are independent standard normal random variables that are the entries of the random vector $\underline{\boldsymbol{\xi}}=U^{T}\underline{\boldsymbol{x}}$. Then
$\mathbb{E}[e^{s\boldsymbol{\Delta}}]=\prod_{i=1}^{n}e^{-s\lambda_{i}}\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}].$ (31)
Using $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $12s\lambda_{i}<$ 1, we get
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{-\frac{1}{2}\ln(1-2s\lambda_{i})},$ (32)
for all $s\geq 0$ such that $12s\lambda_{i}<$ 1 for all $1\leq i\leq n$. Now, let $a,b>0$ be such that (6) holds. Then the bound in (32) can be loosened as
$\Pr(\boldsymbol{\Delta}>t)$$\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+a(2s\lambda_{i})^{2})}$ (33)
$=$ $\exp\Big{(}-st+2as^{2}\sum_{i=1}^{n}\lambda_{i}^{2}\Big{)}$
$=$ $\exp\big{(}-st+2a\|A\|_{2}^{2}s^{2}\big{)},$
for all $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The minimum for the function
$f(s)=-st+2a\|A\|_{2}^{2}s^{2}$
over $0\leq s\leq\frac{b}{2\|A\|}$ occurs at $s_{opt}=\min\{s_{0},\frac{b}{2\|A\|}\}$ where $s_{0}=\frac{t}{4a\|A\|_{2}^{2}}$. We consider two cases:
  1. 1.
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then
    $f(s_{opt})$ $=$ $f(s_{0})$ (34)
    $=$ $-\frac{t^{2}}{4a\|A\|_{2}^{2}}+2a\|A\|_{2}^{2}\Big{(}\frac{t}{4a\|A\|_{2}^{2}}\Big{)}^{2}$
    $=$ $-\frac{t^{2}}{8a\|A\|_{2}^{2}}.$
  2. 2.
    Let $s_{0}>\frac{b}{2\|A\|}$. Then
    $f(s_{opt})$ $=$ $f\Big{(}\frac{b}{2\|A\|}\Big{)}$ (35)
    $=$ $-\frac{bt}{2\|A\|}+2a\|A\|_{2}^{2}\Big{(}\frac{b}{2\|A\|}\Big{)}^{2}$
    $\stackrel{{\scriptstyle}}{{<}}$ $-\frac{bt}{2\|A\|}+2a\|A\|_{2}^{2}\times s_{0}\times\frac{b}{2\|A\|}$
    $=$ $-\frac{bt}{2\|A\|}+\frac{t}{2}\times\frac{b}{2\|A\|}$
    $=$ $-\frac{bt}{4\|A\|}.$
By (34) and (35),
$f(s_{opt})$$\leq$ $\max\Big{\{}-\frac{t^{2}}{8a\|A\|_{2}^{2}},-\frac{bt}{4\|A\|}\Big{\}}$ (36)
$=$ $-\min\Big{\{}\frac{t^{2}}{8a\|A\|_{2}^{2}},\frac{bt}{4\|A\|}\Big{\}}$
$\leq$ $-\kappa(a,b)\min\Big{\{}\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\Big{\}},$
where
$\kappa(a,b)=\min\Big{\{}\frac{1}{8a},\frac{b}{4}\Big{\}}.$ (37)
Next, let us explore the constants $a,b>0$. Reference [3] provides $a=1$ and $b=\frac{1}{2}$. This results in $\kappa=\kappa(1,0.5)=0.125$. But, one can do better. We need the special case of the following lemma for $m=1$:
Lemma 1.
Let $m\geq 1$ be an integer. For every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (6) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$ defined in (7). Equivalently, for every $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (6) holds is $a=\theta_{m}(b)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (6) holds.
Proof.
See Appendix A. ∎
Using Lemma 1 with $m=1$ results in
$\kappa=\kappa(\theta_{1}(b),b)=\min\Big{\{}\frac{1}{8\theta_{1}(b)},\frac{b}{4}\Big{\}}.$ (38)
for arbitrary $10<b<$ 1. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $12b\theta_{1}(b)=$ 1. This is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 1
Figure 1: The maximum for $\kappa(\theta_{1}(b),b)=\min\Big{\{}\frac{1}{8\theta_{1}(b)},\frac{b}{4}\Big{\}}$ is achieved where $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$. This is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$.
IV Proof of Proposition 2
Since $A$ is positive-semidefinite, its eigenvalues are all nonnegative. We begin with the inequality in (32). Assume (10) holds for some $a,b>0$ such that $ab<1$. Then
$\Pr(\boldsymbol{\Delta}>t)$$\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+\frac{(2s\lambda_{i})^{2}}{2(1-2as\lambda_{i})})}$ (39)
$=$ $e^{-st}\prod_{i=1}^{n}e^{\frac{s^{2}\lambda^{2}_{i}}{1-2as\lambda_{i}}}$
$=$ $\exp\Big{(}-st+\sum_{i=1}^{n}\frac{s^{2}\lambda_{i}^{2}}{1-2as\lambda_{i}}\Big{)},$
for all $s\geq 0$ such that $0\leq 2s\lambda_{i}\leq b$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. Since $ab<1$ by assumption, then
$0<1-ab\leq 1-2as\lambda_{i},$ (40)
for every $1\leq i\leq n$. Therefore, $0<1-2as\|A\|\leq 1-2as\lambda_{i}$ for every $1\leq i\leq n$ and we can loosen the upper bound in (39) as
$\Pr(\boldsymbol{\Delta}>t)$$\leq$ $\exp\Big{(}-st+\sum_{i=1}^{n}\frac{s^{2}\lambda_{i}^{2}}{1-2as\|A\|}\Big{)}$ (41)
[…]
At this point, we adopt the notations
$\alpha=\|A\|,\,\,\,\,\,\beta=\|A\|_{2}^{2},$ (42)
for notational simplicity. Let us call the exponent on the right side of (41) by $f(s)$, i.e.,
$f(s)=-st+\frac{\beta s^{2}}{1-2a\alpha s}.$ (43)
It is straightforward to check that the function $f$ is convex and achieves its minimum over its domain $[0,\frac{1}{2a\alpha})$ at
$s_{0}=\frac{1}{2a\alpha}\Bigg{(}1-\frac{1}{\big{(}1+\frac{2a\alpha t}{\beta}\big{)}^{\frac{1}{2}}}\Bigg{)}.$ (44)
It follows that the expression on the right side of (41) achieves its minimum at
$s_{opt}=\min\Big{\{}s_{0},\frac{b}{2\alpha}\Big{\}}.$ (45)
We consider two cases:
  1. 1.
    Let $s_{0}\leq\frac{b}{2\alpha}$, or equivalently,
    $\frac{\alpha t}{\beta}\leq\frac{1}{2a}\Big{(}\frac{1}{(1-ab)^{2}}-1\Big{)}.$ (46)
    Then
    $f(s_{opt})=f(s_{0})=-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\bigg{(}\Big{(}1+\frac{2a\alpha t}{\beta}\Big{)}^{\frac{1}{2}}-1\bigg{)},$ (47)
    where we omit the simple algebra.
  2. 2.
    Let $s_{0}>\frac{b}{2\alpha}$. Then
    $f(s_{opt})=f\Big{(}\frac{b}{2\|A\|}\Big{)}=-\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}$ (48)
Let us denote the expression on the right side of (46) by $c$, i.e.,
$c=\frac{1}{2a}\Big{(}\frac{1}{(1-ab)^{2}}-1\Big{)}.$ (49)
To recap, we have shown that
$\Pr(\boldsymbol{\Delta}>t)\leq e^{f(s_{opt})}$ (50)
and
$f(s_{opt})=\left\{\begin{array}[]{cc}-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\Big{(}\big{(}1+\frac{2a\alpha t}{\beta}\big{)}^{\frac{1}{2}}-1\Big{)}&\frac{\alpha t}{\beta}\leq c\\ -\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}&\frac{\alpha t}{\beta}>c\end{array}\right..$ (53)
Next, the next lemma explores a possibility for $a,b$. It addresses an inequality for which (10) is a special case.
Lemma 2.
Let $m\geq 1$ be an integer. For every $\frac{m+1}{m+2}<a<1$, the inequality
[…]
holds with
$b=\frac{m+1-(m+2)a}{a(m-(m+1)a)}.$ (55)
Proof.
See Appendix B. ∎
Using Lemma 2 for $m=1$, we obtain
$b=\frac{3a-2}{a(2a-1)},\,\,\,\,\,\frac{2}{3}<a<1.$ (56)
Note that $ab=\frac{3a-2}{2a-1}<1$ as required in (40). Moreover, $1-ab=\frac{1-a}{2a-1}$ and the threshold $c$ in (49) is given by
$c=\frac{3a-2}{2(1-a)^{2}}.$ (57)
V Proof of Proposition 3
To see how (13) implies HW inequality, let us write HW inequality in (2) as
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}}.$ (58)
Note
$\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}=\left\{\begin{array}[]{cc}\frac{t^{2}}{\beta}&\frac{\alpha t}{\beta}\leq 1\\ \frac{t}{\alpha}&\frac{\alpha t}{\beta}>1\end{array}\right..$ (61)
We aim to show that there is $\kappa>0$ such that the upper bound in (50) is less than or equal to the upper bound in (58), or equivalently, $f(s_{opt})\leq-\kappa\min\{\frac{\beta}{t^{2}},\frac{\alpha}{t}\}$. We want this to hold regardless of $t,\alpha,\beta$. Thus, we look for largest $\kappa>0$ such that
$\kappa\leq\inf_{t,\alpha,\beta>0}-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (62)
Interestingly, the ratio $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio
$\rho=\frac{\alpha t}{\beta}.$ (63)
For ease of comparing the right sides in (16) and (61), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible domain $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
$a_{0}=\frac{7-\sqrt{17}}{4}\approx 0.7192.$ (64)
One can easily check that the corresponding value for $b$ by $\frac{3a_{0}-2}{a_{0}(2a_{0}-1)}=\frac{1}{2}$. Then we have
$-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a_{0}\rho}-\frac{1}{2a_{0}^{2}\rho^{2}}\big{(}\sqrt{1+2a_{0}\rho}-1\big{)}&\rho\leq 1\\ \frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})\rho}&\rho>1\end{array}\right..$ (67)
Fig. 2 shows the graph of the function of $\rho$ on the right side of (67). It achieves its absolute minimum value at $\rho=1$. It follows that the value for the infimum on the right side of (62) is
$\frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})}=\frac{9-\sqrt{17}}{32}\approx 0.152.$ (68)
Figure 2: Graph of the function of $\rho$ on the right side of (67). It achieves its absolute minimum value at $\rho=1$.
It is shown in Appendix C that the choice of $a=a_{0}$ in (64) results in the largest value for $\kappa$.
Finally, we address LM inequality itself. Let us write LM inequality in (4) slightly differently as follows. Call $2\sqrt{\beta t}+2\alpha t=t^{\prime}$ where $\alpha,\beta$ are as in (42). Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LM inequality becomes
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\left(-\Big{(}\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha}\Big{)}^{2}\right).$ (69)
Consider HW inequality in (58). LM inequality implies HW inequality for all $\kappa$ satisfying
$\kappa\leq\inf_{t,\alpha,\beta>0}-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (70)
Once again, we realize that the ratio on the right side of (70) depends entirely on the ratio $\rho=\frac{\alpha t}{\beta}$. It is given by
$-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}(\frac{\sqrt{1+2\rho}-1}{2\rho})^{2}&\rho\leq 1\\ \frac{(\sqrt{1+2\rho}-1)^{2}}{4\rho}&\rho>1\end{array}\right..$ (73)
One can easily check that this function of $\rho$ achieves it absolute minimum value of $1-\frac{\sqrt{3}}{2}$ at $\rho=1$. This establishes Proposition 3.
VI Proof of Proposition 4
We begin with the bound in (32). By Lemma 1 in Section III, for every integer $m\geq 1$ and $10<b<$ 1, the inequality in (19) is the tightest for $a=\theta_{m}(b)$. Then
$\Pr(\boldsymbol{\Delta}>t)$$\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+a|2s\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{(2s\lambda_{i})^{k}}{k})}$ (74)
$=$ $\exp\Big{(}-st+a2^{m}s^{m+1}\sum_{i=1}^{n}|\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{1}{k}2^{k-1}s^{k}\sum_{i=1}^{n}\lambda_{i}^{k}\Big{)}$
$=$ $\exp\Big{(}-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}+\sum_{k=2}^{m}\frac{2^{k-1}s^{k}\mathrm{Tr}(A^{k})}{k}\Big{)},$
for all $s\geq 0$ such that $2s\lambda_{i}\leq b$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (74) in terms of $s$ is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
$s_{0}=\frac{1}{2}\left(\frac{t}{a(m+1)\|A\|_{m+1}^{m+1}}\right)^{\frac{1}{m}}.$ (75)
We plug in $s=s_{*}$ on the right side in (74) where333
$s_{*}=\min\Big{\{}s_{0},\frac{b}{2\|A\|}\Big{\}}.$ (76)
The third term in the exponent in (74) is bounded by
$\sum_{k=2}^{m}\frac{2^{k-1}s_{*}^{k}\mathrm{Tr}(A^{k})}{k}$$\leq$ $\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}\mathrm{Tr}(A^{k})}{k}$ (77)
$=$ $\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{\mathrm{Tr}(A^{k})}{\|A\|^{k}}$
$\leq$ $\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$
where the last step is due to $\mathrm{Tr}(A^{k})\leq n\|A^{k}\|\leq n\|A\|^{k}$. Next, we bound the sum of the first two terms, namely, $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases:
  1. 1.
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then $s_{*}=s_{0}$ and simple algebra shows that
    $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}=-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}}.$ (78)
    where
    $\kappa_{m,1}=\frac{1-\frac{1}{m+1}}{2a^{\frac{1}{m}}(m+1)^{\frac{1}{m}}}$ (79)
  2. 2.
    Let $s_{0}>\frac{b}{2\|A\|}$. Then $s_{opt}=\frac{b}{2\|A\|}$ and we have
    $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}$ $=$ $-s_{*}t+a2^{m}s_{*}\times s_{*}^{m}\|A\|_{m+1}^{m+1}$ (80)
    $\leq$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times s_{0}^{m}\|A\|_{m+1}^{m+1}$
    $=$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times\frac{t}{a2^{m}(m+1)\|A\|_{m+1}^{m+1}}\|A\|_{m+1}^{m+1}$
    $=$ $-\kappa_{m,2}\frac{t}{\|A\|},$
    where
    $\kappa_{m,2}=\frac{b}{2}\Big{(}1-\frac{1}{m+1}\Big{)}.$ (81)
By (78) and (80),
$-s_{opt}t+a2^{m}s_{opt}^{m+1}\mathrm{tr}(A^{m+1})$$\leq$ $\max\Big{\{}-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},-\kappa_{m,2}\frac{t}{\|A\|}\Big{\}}$ (82)
$=$ $-\min\Big{\{}\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\kappa_{m,2}(a,b)\frac{t}{\|A\|}\Big{\}}$
$\leq$ $-\min\{\kappa_{m,1},\kappa_{m,2}\}\min\bigg{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg{\}}.$
VII Proof of Proposition 5
Consider the bound in (74) for $m=2$, i.e.,
$\Pr(\boldsymbol{\Delta}>t)\leq e^{f(s)},$ (83)
where $f(s)$ is given by
[…]
and $\alpha,\beta$ and $\gamma$ are defined in (22). The function $f(s)$ achieves it global minimum value over $s\geq 0$ at $s=s_{0}$ given by
$s_{0}=\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{12a\gamma}.$ (85)
It follows that the minimum value for $f(s)$ over the interval $0\leq s\leq\frac{b}{2\alpha}$ is achieved at
$s_{opt}=\min\Big{\{}s_{0},\frac{b}{2\alpha}\Big{\}}.$ (86)
Two cases are considered:
  1. 1.
    Let $s_{0}\leq\frac{b}{2\alpha}$. Then $s_{opt}=s_{0}$ and $f(s_{opt})=-s_{0}t+\beta s_{0}^{2}+4a\gamma s_{0}^{3}$. To continue, we need the following lemma.
    Lemma 3.
    Let $u,v,w,w^{\prime}$ be positive numbers such that $w\geq w^{\prime}$ and $x=\frac{\sqrt{v^{2}+3uw}-v}{3w}$. Then
    $-ux+vx^{2}+wx^{3}\leq\frac{1}{27w^{\frac{1}{2}}w^{\prime\frac{3}{2}}}(\sqrt{v^{2}+3uw^{\prime}}-v)(v\sqrt{v^{2}+3uw^{\prime}}-v^{2}-6uw^{\prime}).$ (87)
    Proof.
    See Appendix D. ∎
    Applying Lemma 3 with $x=s_{0}$, $u=t$, $v=\beta$ and $w=4a\gamma$ and noting that $w\geq w^{\prime}=\frac{4}{3}\gamma$ due to $a=\theta_{2}(b)\geq\frac{1}{3}$, we obtain
    $f(s_{opt})\leq\frac{\sqrt{3}}{144a^{\frac{1}{2}}\gamma^{2}}(\sqrt{\beta^{2}+4\gamma t}-\beta)(\beta\sqrt{\beta^{2}+4\gamma t}-\beta^{2}-8\gamma t)$ (88)
  2. 2.
    Let $s_{0}>\frac{b}{2\alpha}$. Then $s_{opt}=\frac{b}{2\alpha}$ and
    $f(s_{opt})$ $=$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma s_{opt}^{3}$ (89)
    $=$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma s_{opt}\times s_{opt}^{2}$
    $\leq$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma\frac{b}{2\alpha}\times s_{0}^{2}.$
    Next, we bound the two terms on the right side of (89). We write the second term as
    $\beta s_{opt}^{2}$ $=$ $\beta s_{opt}\times s_{opt}$ (90)
    $\leq$ $\beta\frac{b}{2\alpha}\times s_{0}$
    $=$ $\frac{\beta b}{24\alpha\gamma}\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{a}$
    $\leq$ $\frac{\beta b}{24\alpha\gamma}\frac{\sqrt{\beta^{2}+12\times\frac{1}{3}\times\gamma t}-\beta}{\frac{1}{3}}$
    $=$ $\frac{\beta b}{8\alpha\gamma}(\sqrt{\beta^{2}+4\gamma t}-\beta),$
    where the penultimate step uses the fact that $a\mapsto\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{a}$ is a decreasing function in terms of $a$ and $a=\theta_{2}(b)\geq\frac{1}{3}$. Alternatively, we can bound $\beta s_{opt}^{2}$ as
    $\beta s_{opt}^{2}=\frac{\beta b^{2}}{4\alpha^{2}}\leq\frac{\beta b}{4\alpha^{2}},$ (91)
    where we simply use the fact that $b^{2}\leq b$ as $10<b<$ 1. Using the tighter bound in (90) and (91), we get
    $\beta s_{opt}^{2}\leq\frac{\beta b}{4\alpha}\min\Big{\{}\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t-\beta}}{2\gamma}\Big{\}}.$ (92)
    Regarding the third term on the right side of (89), note that $s_{0}\leq\frac{\sqrt{\beta^{2}}+\sqrt{12a\gamma t}-\beta}{12a\gamma t}=\sqrt{\frac{t}{12a\gamma}}$ where we used the trivial inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for $x,y\geq 0$. Then
    $4a\gamma\frac{b}{2\alpha}\times s_{0}^{2}\leq 4a\gamma\frac{b}{2\alpha}\times\frac{t}{12a\gamma}=\frac{bt}{6\alpha}.$ (93)
    By (89), (92) and (93),
    $f(s_{opt})$$\leq$ $-\frac{bt}{2\alpha}+\frac{\beta b}{4\alpha}\min\Big{\{}\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t-\beta}}{2\gamma}\Big{\}}+\frac{bt}{6\alpha}$ (94)
    $=$ $\frac{\beta b}{4\alpha}\min\Big{\{}\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t-\beta}}{2\gamma}\Big{\}}-\frac{bt}{3\alpha}$
Appendix A; Proof of Lemma 1
Define the function $f$ by
$f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+a|x|^{m+1}+\ln(1-x).$ (95)
We study the cases $x>0$ and $x<0$, separately.
  • Let $x>0$. Then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ and we get
    $f^{\prime}(x)$ $=$ $1+x+x^{2}+\cdots+x^{m-1}+(m+1)ax^{m}-\frac{1}{1-x}$ (96)
    $=$ $\frac{(1+x+x^{2}+\cdots+x^{m-1})(1-x)+(m+1)ax^{m}(1-x)-1}{1-x}$
    $=$ $\frac{1-x^{m}+(m+1)ax^{m}(1-x)-1}{1-x}$
    $=$ $\frac{x^{m}((m+1)a-1-(m+1)ax)}{1-x}.$
    The only positive first-order critical number for $f$ is $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, one easily checks that $f^{\prime}(x)<0$ over an open interval $I$ around 00. Since $f(0)=0$, then the Mean Value Theorem (MVT) implies that $f(x)<0$ over $I\cap(0,\infty)$ and hence, (6) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $10<x_{0}<$ 1, $f^{\prime}(x)>0$ over $(0,x_{0})$ and $f^{\prime}(x)<0$ over $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=-\infty$. This tells us that $f$ rises about 00 on right of $x=0$, reaches an absolute maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to $\infty$- ∞ on left of $x=1$. As such, $f$ possesses a unique zero (x-crossing) somewhere over $(x_{0},1)$. We let $b$ be this zero. Then (6) clearly holds for $0\leq x\leq b$. Finally, we show that $b$ is $\theta_{m}^{-1}(a)$, or equivalently, $\theta_{m}(b)=a$. We have $f(b)=0$, i.e.,
    $-\ln(1-b)=b+\frac{b^{2}}{2}+\frac{b^{3}}{3}+\cdots+\frac{b^{m}}{m}+ab^{m+1}.$ (97)
    Recalling the Maclaurin series for $-\ln(1-x)$, we have $-\ln(1-b)=\sum_{j=1}^{\infty}\frac{b^{j}}{j}$. Then (97) can be simplified as
    $\sum_{j=m+1}^{\infty}\frac{b^{j}}{j}=ab^{m+1}$ (98)
    and hence,
    $a=\sum_{j=m+1}^{\infty}\frac{b^{j-m-1}}{j}.$ (99)
    Changing the index $j$ to $j+m+1$, we get $a=\sum_{j=0}^{\infty}\frac{b^{j}}{j+m+1}$, i.e., $a=\theta_{m}(b)$ as promised.
  • Now, let $x<0$. If $m$ is odd, $f(x)$ is given by $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ as in the previous case and one easily checks that for $a>\frac{1}{n+1}$, $f^{\prime}(x)$ given in (96) is negative for every $x<0$. Then MVT implies that $f(x)>0$ for every $x<0$ as desired. If $m$ is even, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}-ax^{m+1}+\ln(1-x)$. A similar computation as in (96) gives that
    $f^{\prime}(x)=\frac{-x^{m}((m+1)a+1-(m+1)ax)}{1-x}.$ (100)
    Since $m$ is even, once again we have $f^{\prime}(x)<0$ for every $x>0$ and MVT comes in one last time to confirm that $f(x)>0$ for $x<0$.
    VIII Appendix B; Proof of Lemma 2
    Assume $m\geq 1$ is an integer and fix $\frac{m+1}{m+2}<a<1$. Define the function $f$ by
[…]
    One can easily check that
    $f^{\prime}(x)=\frac{x^{m+1}((m+2)a-m-1-a((m+1)a-m)x)}{(m+1)(1-x)(1-ax)^{2}}.$ (102)
    The only positive first-order critical number for $f$ over the interval $(0,1)$ is given by $x_{0}=\frac{m+1-(m+2)a}{a(m-(m+1)a)}$. Moreover, $f^{\prime}(x)>0$ for $0<x<x_{0}$ and $f^{\prime}(x)<0$ for $x_{0}<x<1$. Since $f(0)=0$, then MVT implies that $f(x)>0$ for all $x$ in the interval $(0,x_{0}]$ and hence, (10) holds with $b=x_{0}$.
Appendix C; The choice of $a=a_{0}$ in (16) results in the largest $\kappa$
In order to compare $f(s_{opt})$ in (53) with $\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}$, we address the cases $c\leq 1$ and $c>1$ separately.
  1. 1.
    Let $c\leq 1$, or equivalently, $\frac{2}{3}<a\leq a_{0}$. Then
    $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq c\\ \frac{b}{2\rho}-\frac{b^{2}}{4(1-ab)\rho^{2}}&c<\rho\leq 1\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>1\end{array}\right..$ (106)
    One can easily check that this is a continuous function of $\rho$ and it achieves its absolute minimum value of $\frac{b}{2}-\frac{b^{2}}{4(1-ab)}=\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}$ at $\rho=1$.
  2. 2.
    Let $c>1$, or equivalently, $a_{0}<a<1$. Then
    $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq 1\\ \frac{1}{2a}-\frac{1}{2a^{2}\rho}(\sqrt{1+2a\rho}-1)&1<\rho\leq c\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>c\end{array}\right..$ (110)
    It is routine to check that this is a continuous function of $a$ and it achieves its absolute minimum value of $\frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)$ at $\rho=1$.
We have shown that
$\min_{\rho>0}-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}&\frac{2}{3}<a\leq a_{0}\\ \frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)&a_{0}<a<1\end{array}\right..$ (113)
We need to choose $\frac{2}{3}<a<1$ such that the right side of (113) is as large as possible. Fig. LABEL:pic_1111 shows the graph of this function over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
Figure 3: The graph of the function of $a$ on the right side of (113) over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$
Appendix D; Proof of Lemma 3
Define the function
$g(u,v,w)=\frac{1}{w^{\frac{3}{2}}}(\sqrt{v^{2}+3uw}-v)(v\sqrt{v^{2}+3uw}-v^{2}-6uw).$ (114)
Tedious algebra shows that
$-ux+vx^{2}+wx^{3}=\frac{g(u,v,w)}{27w^{\frac{1}{2}}}.$ (115)
We show that $g$ is decreasing in $w$ for given $u,v$. Another round of tedious algebra shows
$\frac{\partial g}{\partial w}=-\frac{3v}{2w^{\frac{5}{2}}}\big{(}2v^{2}+3uw-2v\sqrt{v^{2}+3uw}\big{)}.$ (116)
Then $\frac{\partial g}{\partial w}\leq 0$ follows by noting that $2v^{2}+3uw\geq 2v\sqrt{v^{2}+3uw}$ for all $u,v,w$.