### v3

# Refining Concentration for Gaussian Quadratic Chaos

Abstract
We start by slightly modifying the proof of Hanson-Wright inequality (HWI) for concentration of Gaussian quadratic chaos where we tighten the bound by increasing the absolute constant in its formulation from the largest known value of 0.125 to at least 0.145 in the symmetric case. We also present a sharper version of an inequality due to Laurent and Massart (LMI) through which we increase the absolute constant in HWI from the largest available value of approximately $0.134$ due to LMI itself to at least $0.152$ in the positive-semidefinite case. A new sequence of concentration bounds indexed by $m=1,2,3,\cdots,\infty$ is developed that involves Schatten norms of the underlying matrix. The case $m=1$ recovers HWI. These bounds undergo a phase transition in the sense that if the tail parameter is smaller than a critical threshold $\tau_{c}$, then $m=1$ is the tightest and if it is larger than $\tau_{c}$, then $m=\infty$ is the tightest. This leads to a novel and simple bound referred to as the $m_{\infty}$-bound that depends on the underlying matrix only through its operator norm. A separate concentration bound named “twin to HWI” is also developed that is tighter than HWI for both sufficiently small and large values of the tail parameter. The twin bound involves Schatten norms of orders $2$ and $3$ and the operator norm. To further investigate the $m_{\infty}$-bound, we explore concentration bounds when the underlying matrix is positive-semidefinite and only the dimension $n$ and its largest eigenvalue are known. Five candidates are examined, namely, the $m_{\infty}$-bound, relaxed versions of HWI and LMI, the $\chi^{2}$-bound and the large deviations bound. The sharpest among these is always either the $m_{\infty}$-bound or the $\chi^{2}$-bound. The case of even dimension is given special attention. If $n=2,4,6$, the $\chi^{2}$-bound is tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, the $m_{\infty}$-bound is sharper than the $\chi^{2}$-bound if and only if the ratio of the tail parameter over the largest eigenvalue lies inside a finite open interval which expands indefinitely as $n$ grows.
I Introduction
I-A Summary of prior art
This paper concerns the Gaussian version of Hanson-Wright inequality (HWI) [1], a concentration of measure result that presents an upper bound on the tail probability for a quadratic form in a random vector of independent standard Gaussian random variables also known as order 2 chaos. [^1] In its most general form, HWI concerns quadratic forms in sub-Gaussian random vectors. For an integer $n\geq 1$ called the dimension throughout the paper, let $\underline{\boldsymbol{x}}$ be such a random vector of length $n$ and $A$ be an arbitrary $n\times n$ matrix. Denote the difference between $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}$ and its expected value $\mathbb{E}[\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}]=\mathrm{Tr}(A)$ by $\boldsymbol{\Delta}$, i.e.,
$\boldsymbol{\Delta}=\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}-\mathrm{Tr}(A).$ (1)
HWI states that for every tail parameter $t>0$,
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\Big(-\kappa\min\Big\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\Big\}\Big),$ (2)
where $\|A\|_{2}$ is the Hilbert-Schmidt norm of $A$, $\|A\|$ is the operator norm of $A$ and $\kappa$ is an absolute constant that does not depend on $n$, AA and tt. [^2] Replacing AA by A-A in (2) results in the same upper bound as in (2) on the left-tail probability $\Pr(\boldsymbol{\Delta}<-t)$. Without loss of generality, we assume $A$ is symmetric throughout the paper. This can be easily justified by noting that $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}=\underline{\boldsymbol{x}}^{T}\tilde{A}\underline{\boldsymbol{x}}$ where $\tilde{A}=\frac{1}{2}(A+A^{T})$ is symmetric and $\|\tilde{A}\|_{2}\leq\|A\|_{2}$ and $\|\tilde{A}\|\leq\|A\|$. These inequalities follow by applying the triangle inequality for norms and noting that $\|A\|_{2}=\|A^{T}\|_{2}$ and $\|A\|=\|A^{T}\|$. In the original paper [1], $\|A\|$ is the operator norm of the matrix whose entries are absolute values of entries of $A$. Reference [2] shows that the inequality indeed holds as stated in (2) with $\|A\|$ being the operator norm of $A$ itself. An explicit value for $\kappa$ is missing or at least hard to identify in [1, 2]. Theorem 3.1.4 on page 75 in [3] provides the bounds
$\Pr(\boldsymbol{\Delta}>t)\leq\left\{\begin{array}[]{cc}\exp(-\frac{t^{2}}{8\|A\|_{2}^{2}})&0<t\leq\frac{\|A\|_{2}^{2}}{\|A\|}\\ \exp(\frac{\|A\|_{2}^{2}}{8\|A\|^{2}}-\frac{t}{4\|A\|})&t>\frac{\|A\|_{2}^{2}}{\|A\|}\end{array}\right.$. (5)
For $t>\frac{\|A\|_{2}^{2}}{\|A\|}$, we have $\frac{\|A\|_{2}^{2}}{8\|A\|^{2}}-\frac{t}{4\|A\|}<-\frac{t}{8\|A\|}$. Therefore, (5) implies HWI with $\kappa=\frac{1}{8}=0.125$. More recently, Textbook [4] gives the explicit value $\kappa=\frac{1}{8}$ at the end of the proof for Theorem B.8 on page 304. This value for $\kappa$ can also be inferred form inequality (3.28) on page 120 in [5] given by
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\Big(-\frac{t^{2}}{4(\|A\|_{2}^{2}+\|A\|t)}\Big).$ (6)
It is stated in [5] that (6) holds for arbitrary symmetric matrix $A$. However, there is a mistake in the proof. Inequality $\frac{1}{2}(-\ln(1-2x)-2x)\leq\frac{x^{2}}{1-2x}$ in the middle of page 120 does not hold for $x<0$. Nonetheless, inequality (6) certainly holds when $A$ is positive-semidefinite. Loosening the bound by writing $\|A\|_{2}^{2}+\|A\|t\leq 2\max\{\|A\|_{2}^{2},\|A\|t\}$, we arrive at HWI with $\kappa=\frac{1}{8}$. For a positive-semidefinite matrix $A$, a key result in the literature on concentration of Gaussian quadratic chaos is inequality (4.1) on page 1325 in [6] due to Laurent and Massart given by
$\Pr(\boldsymbol{\Delta}>2\|A\|_{2}\sqrt{t}+2\|A\|t)\leq\exp(-t),$ (7)
for all tail parameters $t>0$. We will refer to (7) as Laurent-Massart inequality (LMI). It is notable that LMI also appears in Example 2.12 in Textbook [7]. One can show that LMI is equivalent to [^3] See Section V.
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\bigg(-\Big(\frac{\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2}}{2\|A\|}\Big)^{2}\bigg).$ (8)
We will show in Section V that LMI implies HWI at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
I-B Contributions and main results
The main contributions in this paper are fivefold.
  1. 1.
    We visit and slightly modify the proof given in [4] for HWI. In the course of proof, Reference [4] uses the bound
    $-\ln(1-x)\leq x+x^{2},\,\,\,\,\,|x|\leq\frac{1}{2},$ (9)
    which ultimately results in $\kappa=0.125$. Instead, we consider
    $-\ln(1-x)\leq x+ax^{2},\,\,\,\,\,|x|\leq b,$ (10)
    for $a,b>0$. For given $0<b<1$, we determine the smallest $a$ (tightest bound) such that (10) holds. As a result, we are able to increase $\kappa$ from $0.125$ to at least $0.145$ for arbitrary symmetric matrix $A$. More precisely, we have the following proposition.
    Proposition 1.
    For $0<b<1$ and integer $m\geq 1$ define the sequence of functions
    $\theta_{m}(b)=-\frac{1}{b^{m+1}}\Big(\ln(1-b)+b+\frac{b^{2}}{2}+\frac{b^{3}}{3}+\cdots+\frac{b^{m}}{m}\Big).$ (11)
    If $A$ is symmetric, then HWI holds with
    $\kappa=\frac{b^{*}}{4}\approx 0.1457,$ (12)
    where $b^{*}$ is the unique solution for $b$ in the equation $2b\theta_{1}(b)=1$.
    Proof.
    See Section III.∎
    A special case of Lemma 1 in Section III shows that the smallest value for the parameter $a$ in (10) is $\theta_{1}(b)$. This optimal value for $a$ is also reported independently in Reference [8] in Corollary 4 on Page 5. The functions $\theta_{m}(b)$ will appear again in Theorem 1 and Theorem 2 in below.
  2. 2.
    In the positive-semidefinite case, we present a sharper version of LMI. To prove (7), Reference [6] relies on the bound
    $-\ln(1-x)\leq x+\frac{x^{2}}{2(1-x)},\,\,\,\,\,0\leq x\leq 1.$ (13)
    Instead, we consider
    $-\ln(1-x)\leq x+\frac{x^{2}}{2(1-ax)},\,\,\,\,\,0\leq x\leq b,$ (14)
    for $a,b>0$. We show that for every $\frac{2}{3}<a\leq 1$, this inequality holds for $b$ at least as large as $\frac{3a-2}{a(2a-1)}$. As a result, we derive an improved LMI as presented in the next proposition.
    Proposition 2.
    Let $A$ be positive-semidefinite and
    $\alpha=\|A\|,\,\,\,\,\,\beta=\|A\|_{2}^{2}.$ (15)
    For $\frac{2}{3}<a\leq 1$, let  [^4] It is understood that $c=\infty$ when $a=1$.
    $b=\frac{3a-2}{a(2a-1)},\,\,\,\,\,c=\frac{3a-2}{2(1-a)^{2}}.$ (16)
    Then
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda(t,a)},$ (17)
    where
    $\Lambda(t,a)=\left\{\begin{array}[]{cc}-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\Big(\sqrt{1+\frac{2a\alpha t}{\beta}}-1\Big)&\frac{\alpha t}{\beta}\leq c\\ -\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}&\frac{\alpha t}{\beta}>c\end{array}\right.$. (20)
    Proof.
    See Section IV. ∎
    Choosing $a=b=1$ and $c=\infty$ in (17) recovers LMI in (8). In Subsection II-A, we minimize the upper bound on the right side of (17) in terms of $a$ and compare the resulting optimized inequality with LMI to demonstrate the improvement. We will see that the minimizing value for $a$ is the unique root inside the interval $(\frac{2}{3},1)$ of a quintic polynomial equation [^5] See Equation (46). and it depends on the matrix $A$ and the tail parameter $t$ only through the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$. We are able to approximate this unique root by an analytic formula [^6] See Equation (41). where the approximation error is shown to be less than $0.035$ regardless of $A$ and $t$. As a consequence of Proposition 2, we offer the next result which is a stronger version of Proposition 1 in the positive-semidefinite case.
    Proposition 3.
    Let $A$ be positive-semidefinite. The inequality in (17) implies HWI at best with
    $\kappa=\frac{9-\sqrt{17}}{32}\approx 0.1524,$ (21)
    which is achieved for
    $a=\frac{7-\sqrt{17}}{4}.$ (22)
    Moreover, LMI in (7) implies HWI at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
    Proof.
    See Section V. ∎
  3. 3.
    We explore beyond HWI for arbitrary symmetric matrix $A$ by considering an extension of (10), i.e., an inequality of the form
    $-\ln(1-x)\leq x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+a|x|^{m+1},\,\,\,\,\,|x|\leq b,$ (23)
    for a given integer $m\geq 1$ and $a,b>0$. We show [^7] See Lemma 1 in Section III. that for every $m\geq 1$ and $0<b<1$, the smallest $a$ (tightest upper bound) for which (23) holds is $\theta_{m}(b)$ as defined in (11). Consequently, we obtain a sequence of concentration bounds for Gaussian quadratic chaos that are written in terms of Schatten norms of AA. [^8] See Subsection I.C for the definition of Schatten $p$-norm.
    Theorem 1.
    Let $m\geq 1$ be an integer, $0<b<1$ and $\|A\|_{m+1}$ be Schatten norm of order $m+1$ for a symmetric matrix $A$. Then
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda_{m}(t,b)},$ (24)
    where [^9] The function $\Lambda(t,a)$ in (20) and the functions $\Lambda_{m}(t,b)$ in (25) are not related.
    $\Lambda_{m}(t,b)=\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k}-\kappa_{m}(b)\min\bigg\{\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg\}$ (25)
    and
    $\kappa_{m}(b)=\frac{m}{2(m+1)}\min\bigg\{\frac{1}{\big((m+1)\theta_{m}(b)\big)^{\frac{1}{m}}},b\bigg\}$ (26)
    with $\theta_{m}(b)$ as defined in (11).
    Proof.
    See Section VI. ∎
    Choosing $m=1$ in (24) recovers HWI in (2) with $\kappa=\kappa_{1}(b)$. Schatten norm $\|A\|_{m+1}$ should not be mistaken with the $L_{p}$-norm of $A$ with $p=m+1$ unless $m=1$ in which case $\|A\|_{2}$ is Hilbert-Schmidt norm of $A$. The upper bound $e^{\Lambda_{m}(t,b)}$ in (24) is to be minimized over $0<b<1$ and the choice of $m\geq 1$. In general, the sequence of upper bounds $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for $m=1,2,3,\cdots$ is not monotone in $m$ for given $A$ and $t$. However, we conjecture that it undergoes a phase transition as stated next.
    Conjecture 1.
    For a given symmetric matrix $A$, there exists a critical threshold $\tau_{c}>0$ such that if $t<\tau_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest for $m=1$ and if $t>\tau_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. [^10] More precisely, for $t>\tau_{c}$, the limit $\lim_{m\to\infty}\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ exists and it is smaller than $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for every $m\geq 1$.
    If Conjecture 1 is true, then one must always choose $m=1$ or let $m$ grow to infinity. In Subsection II-B, we will study the bounds in (24) both through numerical examples and analytically where two corollaries to Theorem 1 are stated. The first corollary looks into the limiting value of (24) as $m$ grows large. This leads to the simple bound
    $\Pr(\boldsymbol{\Delta}>t)\leq\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}},$ (27)
    which we refer to as the $m_{\infty}$-bound or $m_{\infty}$-inequality. The right hand side in (27) depends on $A$ only through the dimension $n$ and its operator norm $\|A\|$. The second corollary presents an estimate (upper bound) for the critical threshold $\tau_{c}$ mentioned in Conjecture 1 in above.
  4. 4.
    The upper bounds in Theorem 1 become tighter than the upper bound in HWI only when $t>\tau_{c}$, i.e., for all sufficiently large values of $t$. We raise the question whether one can improve over HWI for all sufficiently small values of $t$. We answer this question in the affirmative. We are able to present what we refer to as a “twin inequality” to HWI which is tighter than HWI not only for all sufficiently small $t$, but also for all sufficiently large $t$. This bound depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$.
    Theorem 2.
    Let $A$ be a symmetric matrix and
    $\alpha=\|A\|,\,\,\,\beta=\|A\|_{2}^{2},\,\,\,\gamma=\|A\|_{3}^{3}.$ (28)
    Also, let
    $\eta_{1}(t)=\frac{1}{12\gamma^{2}}(\sqrt{\beta^{2}+4\gamma t}-\beta)(\beta^{2}+8\gamma t-\beta\sqrt{\beta^{2}+4\gamma t}\,)$ (29)
    and
    $\eta_{2}(t)=\frac{t}{\alpha}-\frac{3\beta}{4\alpha}\min\bigg\{\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t}-\beta}{2\gamma}\bigg\}.$ (30)
    Then
    $\Pr(\boldsymbol{\Delta}>t)\leq\exp\big(-\kappa^{\prime}\min\{\eta_{1}(t),\eta_{2}(t)\}\big),$ (31)
    where $\kappa^{\prime}=\frac{b^{*}}{3}\approx 0.187$ with $b^{*}$ as the unique solution for $b$ in the equation $\frac{4}{\sqrt{3}}b\sqrt{\theta_{2}(b)}=1$ and $\theta_{2}(b)$ as defined in (11).
    Proof.
    See Section VII. ∎
    The bound in (31) is named a twin to HWI in the sense that $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})$ as $t\to 0$ and $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t}{\|A\|}+O(1)$ as $t\to\infty$. These statements are verified in Subsection II-C. Hence, the inequality in (31) is a look-alike to HWI both as $t\to 0$ and $t\to\infty$. Note that the constant $\kappa^{\prime}\approx 0.187$ is slightly larger than the constant $\kappa\approx 0.145$ in HWI. This is exactly why the twin bound is sharper than HWI for both small and large values of $t$.
  5. 5.
    To further explore the $m_{\infty}$-bound, we raise and attempt to answer the following question in Subsection II-D.
    Let $A$ be an $n\times n$ positive-semidefinite matrix for which we only know the dimension $n$ and the value of its operator norm $\|A\|$ denoted by $\nu$. How much is $\sup_{A\in\mathbb{S}^{n}_{+}:\|A\|=\nu}\Pr(\boldsymbol{\Delta}>t)$ in terms of $n,\nu,t$? Here, $\mathbb{S}^{n}_{+}$ is the cone of $n\times n$ positive-semidefinite matrices.
    We present five bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through $n$ and $\|A\|$. One candidate is the $m_{\infty}$-bound in (27). The second and third candidates are relaxed versions of HWI and LMI, respectively, which do not depend on $\|A\|_{2}$ anymore. The relaxed HWI is
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\big\{\frac{t^{2}}{n\|A\|^{2}},\frac{t}{\|A\|}\big\}},\,\,\,\,\,\kappa=\frac{9-\sqrt{17}}{32}$ (32)
    and the relaxed LMI is given by
    $\Pr(\boldsymbol{\Delta}>t)\leq e^{-\frac{n}{4}\big(\sqrt{1+\frac{2t}{n\|A\|}}-1\big)^{2}}.$ (33)
    The fourth candidate is based on the $\chi^{2}$ distribution and is referred to as the $\chi^{2}$-bound. It is given by
    $\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $1-F_{\chi^{2}_{n}}\Big(1+\frac{t}{\|A\|}\Big),$ (34)
    where $F_{\chi^{2}_{n}}(\cdot)$ is the cumulative distribution function (CDF) for a $\chi^{2}$ random variable with $n$ degrees of freedom. The fifth candidate is a loosened version of the $\chi^{2}$-bound known as the large deviations bound given by
    $\Pr(\boldsymbol{\Delta}>t)<\frac{1}{\sqrt{e}}\Big(\frac{e(1+\frac{t}{\|A\|})}{n}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}},\,\,\,\,\,\frac{t}{\|A\|}\geq n-1.$ (35)
    We emphasize that (35) is only valid for $\frac{t}{\|A\|}\geq n-1$. Derivations of the relaxed HWI, the relaxed LMI, the $\chi^{2}$-bound and the large deviations bound are presented in Subsection II-D. Note that all of the bounds in (27), (32), (33), (34) and (35) depend on $A$ and $t$ only through the dimension $n$ and the ratio $\frac{t}{\|A\|}$. Our first result is the following.
    Proposition 4.
    The $m_{\infty}$-bound in (27) is sharper than the relaxed HWI in (32), the relaxed LMI in (33) and the large deviations bound in (35) for every positive-semidefinite matrix $A$ and $t>0$.
    Proof.
    See Section VIII. ∎
[…]
    Proposition 5.
    If $n=2,4,6$, then the $\chi^{2}$-bound is sharper than the $m_{\infty}$-bound regardless of $A$ and $t$. If $n$ is an even integer greater than or equal to $8$, there exist positive constants $r_{n}$ and $r^{\prime}_{n}$ such that $r_{n}<1<r^{\prime}_{n}$ and the $m_{\infty}$-bound is sharper than the $\chi^{2}$-bound if and only if $r_{n}<\frac{t}{\|A\|}<r^{\prime}_{n}$.
    Proof.
    The proof uses Descartes’ rule of signs. See Section IX. ∎
    Additionally, we prove in Appendix E that as the even integer $n$ increases, $r_{n}$ goes to zero and $r^{\prime}_{n}$ grows to infinity, i.e.,
    $\lim_{n\to\infty}r_{n}=0,\,\,\,\,\,\lim_{n\to\infty}r^{\prime}_{n}=\infty,$ (36)
    where it is understood that $n$ grows inside the set of even integers larger than or equal to $8$. As such, supremacy of the $m_{\infty}$-bound over the $\chi^{2}$-bound becomes stronger as $n$ grows.
I-C Notations
Random variables are denoted by boldfaced letters such as $\boldsymbol{x}$ with expectation $\mathbb{E}[\boldsymbol{x}]$. Vectors are denoted by an underline such as $\underline{\boldsymbol{x}}$. The (real) eigenvalues of an $n\times n$ symmetric matrix $A$ are denoted by $\lambda_{1},\cdots,\lambda_{n}$. The operator norm of $A$ is
$\|A\|=\max_{1\leq i\leq n}|\lambda_{i}|$ (37)
and Schatten $p$-norm of $A$ for $p\geq 1$ is defined by
$\|A\|_{p}=\big(\sum_{i=1}^{n}|\lambda_{i}|^{p}\big)^{1/p}.$ (38)
The case $p=2$ reduces to Hilbert-Schmidt norm of $A$. If $A$ is positive-semidefinite, then
$\|A\|_{p}=(\mathrm{Tr}(A^{p}))^{1/p},$ (39)
where $\mathrm{Tr}(\cdot)$ is the trace operator.
II Analysis of the results in Subsection I-B
In this section, we will study several of the results presented in Subsection I-B.
II-A Improving LMI
Consider the improved LMI given in (17) in Proposition 2. It is easily checked that $\Lambda(t,a)<0$ for all $t>0$ and $\frac{2}{3}<a\leq 1$ and hence, the bound in (17) is nontrivial. Let us denote the ratio $\frac{\alpha t}{\beta}$ by $\rho$, i.e.,
$\rho=\frac{\alpha t}{\beta}.$ (40)
Recall $c=\frac{3a-2}{2(1-a)^{2}}$ in (16). The inequalities $\rho>c$ and $\rho\leq c$ given in the formulation of $\Lambda(t,a)$ in (20) are solved for the parameter $a$ as $\frac{2}{3}<a<\hat{a}_{opt}$ and $\hat{a}_{opt}\leq a<1$, respectively, where $\hat{a}_{opt}$ is defined by
$\hat{a}_{opt}=\frac{4\rho+3-\sqrt{8\rho+9}}{4\rho}.$ (41)
The reason for choosing this notation becomes clear in a moment. The value of $\hat{a}_{opt}$ always lies in the admissible interval $(\frac{2}{3},1)$ and one can check that
$\lim_{\rho\to 0^{+}}\hat{a}_{opt}=\frac{2}{3},\,\,\,\,\lim_{\rho\to\infty}\hat{a}_{opt}=1.$ (42)
Replacing $b$ and $c$ by their values in terms of $a$ and after some algebra, we find
$\Lambda(t,a)=\left\{\begin{array}[]{cc}\frac{t}{\alpha}\frac{(3a-2)(2\rho a^{2}+(3-2\rho)a-2)}{4\rho a^{2}(2a-1)(1-a)}&\frac{2}{3}<a<\hat{a}_{opt}\\ \frac{t}{\alpha}\big(-\frac{1}{2a}+\frac{1}{2a^{2}\rho}(\sqrt{1+2a\rho}-1)\big)&\hat{a}_{opt}\leq a<1\end{array}\right.$. (45)
It is easily seen that $\Lambda(t,a)$ is continuous over the whole interval $\frac{2}{3}<a<1$ and it is strictly increasing over the interval $\hat{a}_{opt}\leq a<1$. As such, the absolute minimum value for $\Lambda(t,a)$ in terms of $a$ occurs somewhere over the interval $(\frac{2}{3},\hat{a}_{opt}]$. Looking for the critical numbers of the function $a\mapsto\Lambda(t,a)$ over the interval $(\frac{2}{3},\hat{a}_{opt})$ leads us to the quintic polynomial equation
$12\rho a^{5}+(36-40\rho)a^{4}+(48\rho-99)a^{3}+(104-24\rho)a^{2}+(4\rho-48)a+8=0.$ (46)
It is not hard to show that this equation has a unique root denoted by $a_{opt}$ inside the aforementioned interval which must be solved for numerically. Note that both $a_{opt}$ and $\hat{a}_{opt}$ depend on the matrix $A$ and the tail parameter $t$ only through the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$. Fig. 1 shows the graph for $\hat{a}_{opt}-a_{opt}$ in terms of $0\leq\rho\leq 500$. This numerical observation suggests that $\hat{a}_{opt}-a_{opt}$ never exceeds $0.035$. Having this evidence, we will use $\hat{a}_{opt}$ in place of $a_{opt}$. We will name
$\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda(t,\hat{a}_{opt})},$ (47)
the augmented LMI and
$\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda(t,a_{opt})},$ (48)
the optimal LMI.
Figure 1: The plot for $\hat{a}_{opt}-a_{opt}$ in terms of $\rho$. Both $a_{opt}$ and $\hat{a}_{opt}$ depend on the matrix $A$ and the tail parameter $t$ through the ratio $\rho$. We observe that $\hat{a}_{opt}-a_{opt}<0.035$ regardless of the value of $\rho$.
As mentioned earlier in Section I, the original LMI itself is recovered from (17) by the choices $a=b=1$ and $c=\infty$ and hence, it is given by
$\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda(t,1)}.$ (49)
Once again, using the fact that $\Lambda(t,a)$ as a function of $a$ is strictly increasing over the interval $\hat{a}_{opt},1]$, we conclude that the augmented LM bound in (47) is tighter than the classic LM bound in (49). We demonstrate the improvement through an example. Let $n=3$ and the positive-semidefinite matrix $A$ be given by [^12] The positive-semidefinite matrix in (50) is randomly generated.
$A=\begin{bmatrix}3.2504&-2.0401&1.9337\\ -2.0401&2.0554&0.3603\\ 1.9337&0.3603&4.5310\end{bmatrix}$. (50)
Figure 2: Panel (a) presents plots of the upper bounds in (47), (48) and (49) in terms of $0\leq t\leq 50$ for the positive-semidefinite matrix $A$ given (50). The difference between $e^{\Lambda(t,a_{opt})}$ and $e^{\Lambda(t,\hat{a}_{opt})}$ is not noticeable in this plot. Panel (b) skectches the difference $e^{\Lambda(t,\hat{a}_{opt})}-e^{\Lambda(t,a_{opt})}$.
Panel (a) in Fig. 2 shows the bounds $e^{\Lambda(t,1)},e^{\Lambda(t,a_{opt})}$ and $e^{\Lambda(t,\hat{a}_{opt})}$ in terms of $0\leq t\leq 50$. The difference between $e^{\Lambda(t,a_{opt})}$ and $e^{\Lambda(t,\hat{a}_{opt})}$ is not noticeable in this plot. Panel (b) plots the difference $e^{\Lambda(t,\hat{a}_{opt})}-e^{\Lambda(t,a_{opt})}$.
II-B Beyond HWI
Next, we investigate the sequence of bounds presented in Theorem 1. Recall that the choice $m=1$ recovers the classic HWI. Two initial remarks are in order.
  1. 1.
    As we mentioned earlier, the upper bound $e^{\Lambda_{m}(t,b)}$ must be minimized over $0<b<1$. Such minimization results in meaningful bounds in the sense that $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq 1$. This is due to the simple observation that $\lim_{b\to 0^{+}}\Lambda_{m}(t,b)=0$.
  2. 2.
    The coefficient $\kappa_{m}(b)$ in the exponent achieves a global maximum value somewhere inside the interval $0<b<1$. Since $\lim_{b\to 0^{+}}\theta_{m}(b)=\theta_{m}(0)=\frac{1}{m+1}$ and $\lim_{b\to 1^{-}}\theta_{m}(b)=\infty$, we have $\lim_{b\to 0^{+}}\kappa_{m}(b)=\lim_{b\to 1^{-}}\kappa_{m}(b)=0$. Fig. 3 plots $\kappa_{m}(b)$ in terms of $b$ for several values of $m$.
    Figure 3: Plots of $\kappa_{m}(b)$ in terms of $0<b<1$ for $m=1,2,5,10$.
Next, we will study the bounds $e^{\Lambda_{m}(t,b)}$ for $m\geq 1$ both numerically and analytically. For our numerical study, we let $n=3$ and fix the symmetric matrix $A$ given by [^13] The symmetric matrix in (51) is randomly generated.
$A=\begin{bmatrix}-0.5352&0.1436&-0.2132\\ 0.1436&-2.1746&-0.3521\\ -0.2132&-0.3521&-0.0571\end{bmatrix}$. (51)
This matrix is not positive-semidefinite as it has both positive and negative eigenvalues.
Figure 4: Panel (a) presents plots of the upper bound $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the infimum $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest for $m=1$, i.e., HWI offers the tightest upper bound. Panel (b) presents similar plots for $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value decreases as $m$ grows.
Fig. 4 in panel (a) presents plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the achieved infimum value $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest when $m=1$, i.e., HWI offers the best bound. In contrast, panel (b) shows plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for the larger tail parameter $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value occurs when $m=17$, i.e, when $m$ is the largest. This is an indication of a phase transition in the behaviour of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of the pair of parameters $(m,t)$. As stated in Conjecture 1 in Subsection I-B, there exists a critical threshold $\tau_{c}$ such that if $t<\tau_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m=1$ and if $t>\tau_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. To better observe this phase transition, Fig. 5 offers plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for the values $5\leq t\leq 10$ in steps of $0.2$, i.e., $t=5,5.2,5.4,\cdots,9.8,10$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. We have $m_{opt}=1$ for $t\leq 7$ and $m_{opt}=20$ for $t\geq 7.2$. This tells us that if Conjecture 1 is true, then the critical threshold must satisfy $\tau_{c}\leq 7.2$. In Corollary 2 in below, we will present a theoretical upper bound on $\tau_{c}$.
Figure 5: Panel (a) presents plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for $5\leq t\leq 10$ in steps of $0.2$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. If Conjecture 1 is true, then these plots imply that $\tau_{c}\leq 7.2$.
We continue by stating two corollaries to Theorem 1. For our first corollary, we study the behaviour of the bounds in Theorem 1 as $m$ grows to infinity. This will result in what we refer to as the $m_{\infty}$-inequality.
Corollary 1.
For a symmetric matrix $A$,
$\Pr(\boldsymbol{\Delta}>t)\leq\liminf_{m\to\infty}\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}.$ (52)
Proof.
The first inequality in (52) is trivial. For the second inequality, we note that [^14] If $f_{n}:\mathcal{A}\to\mathbb{R}$ is an arbitrary sequence of functions on a set $\mathcal{A}$ of real numbers, then $\inf_{a\in\mathcal{A}}f_{n}(a)\leq f_{n}(x)$ for every $x\in\mathcal{A}$ and $n\geq 1$. Taking limit inferior from both sides, we get $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)\leq\liminf_{n\to\infty}f_{n}(x)$ for every $x\in\mathcal{A}$. This says $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)$ is a lower bound on the set of numbers $\liminf_{n\to\infty}f_{n}(x)$ for $x\in\mathcal{A}$. Thus, $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)\leq\inf_{x\in\mathcal{A}}\liminf_{n\to\infty}f_{n}(x)$ as the infimum of a set is its largest lower bound.
$\liminf_{m\to\infty}\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq\inf_{0<b<1}\liminf_{m\to\infty}e^{\Lambda_{m}(t,b)},$ (53)
for all $t\geq 0$. The limit $\lim_{m\to\infty}\Lambda_{m}(t,b)$ exists. To see this, first we derive a different formula for $\theta_{m}(b)$. Recall the Maclaurin series $\ln(1-b)=\sum_{i=1}^{\infty}\frac{b^{i}}{i}$. Then
$\theta_{m}(b)=-\frac{1}{b^{m+1}}\Big(-\sum_{i=m+1}^{\infty}\frac{b^{i}}{i}\Big)=\sum_{i=m+1}^{\infty}\frac{b^{i-m-1}}{i}.$ (54)
Changing the index $i$ to $i+m+1$ results in $\theta_{m}(b)=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}$. We have
$\frac{1}{m+1}=\frac{b^{0}}{0+m+1}\leq\theta_{m}(b)\leq\sum_{i=0}^{\infty}\frac{b^{i}}{m+1}=\frac{1}{m+1}\sum_{i=0}^{\infty}b^{i}=\frac{1}{(1-b)(m+1)}.$ (55)
Since $\lim_{m\to\infty}(m+1)^{\frac{1}{m}}=1$, it follows by squeezing that
$\lim_{m\to\infty}(\theta_{m}(b))^{\frac{1}{m}}=1.$ (56)
This verifies
$\lim_{m\to\infty}\kappa_{m}(b)=\frac{1}{2}\min\{1,b\}=\frac{b}{2}.$ (57)
It is also evident that
$\lim_{m\to\infty}\|A\|_{m+1}^{1+\frac{1}{m}}=\|A\|$ (58)
and
$\lim_{m\to\infty}\sum_{k=2}^{m}\frac{b^{k}}{k}=\sum_{k=2}^{\infty}\frac{b^{k}}{k}=-\ln(1-b)-b,$ (59)
where we have used the MacLaurin series for $\ln(1-b)$ one more time. By (57), (58) and (59),
$\lim_{m\to\infty}\Lambda_{m}(t,b)=-\frac{n}{2}(\ln(1-b)+b)-\frac{bt}{2\|A\|}.$ (60)
The right side of (60) is minimized in terms of $b$ for
$b=b_{opt}=\frac{t}{n\|A\|+t}.$ (61)
Plugging this value for $b$ on the right side of (60), we arrive at
$\inf_{0<b<1}\lim_{m\to\infty}e^{\Lambda_{m}(t,b)}=\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}.$ (62)
By (53) and (62), the proof is complete. ∎
Using the inequality $(1+x)^{\frac{1}{x}}<e$ for every $x>0$, the $m_{\infty}$-bound in Corollary 1 is nontrivial in the sense that it is less than one.
Our second corollary enables us to find an upper bound on the critical threshold $\tau_{c}$ in Conjecture 1.
Corollary 2.
Let $\hat{\tau}_{c}$ be the unique positive solution for $t$ in the equation
$\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}\exp\Big(-\frac{(\frac{1}{2}-\kappa)t}{\|A\|}\Big)=1,$ (63)
where $\kappa$ is the constant in HWI given in Proposition 1. If Conjecture 1 is true, then $\tau_{c}\leq\hat{\tau}_{c}$.
Proof.
Define the functions
$f(t)=\exp\Big(-\kappa\min\Big\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\Big\}\Big)$ (64)
and
$g(t)=\liminf_{m\to\infty}\inf_{0<b<1}e^{\Lambda_{m}(t,b)}.$ (65)
By definition, the critical threshold $\tau_{c}$ is given by
$\tau_{c}=\sup\{t>0:g(t)\geq f(t)\}.$ (66)
Clearly, $f(t)\geq e^{-\frac{\kappa t}{\|A\|}}$ and $g(t)\leq(1+\frac{t}{n\|A\|})^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}$ due to Corollary 1. It follows that
$\tau_{c}\leq\hat{\tau}_{c}=\sup\Big\{t>0:\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}\geq e^{-\frac{\kappa t}{\|A\|}}\Big\}.$ (67)
As an application of Corollary 2, let us consider the symmetric matrix $A$ in (51). Using the plots in Fig. 5, we concluded that $\tau_{c}\leq 7.2$. Using Corollary 2, we get the sharper bound $\tau_{c}\leq\hat{\tau}_{c}=6.191$.
II-C Improvement over HWI for both all sufficiently small $t$ and large $t$
Here, we look at the twin inequality to HWI in Theorem 2. For a numerical study, let us consider the symmetric matrix $A$ in (51). Panels (a) and (b) in Fig. 6 compare the bound in (31) with HWI for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We observe that inequality (31) is tighter than HWI for both sufficiently small and large values of the tail parameter $t$. As briefly pointed out in Section I, this advantage can be verified by writing asymptotic expansions for the two functions $\eta_{1}(t)$ and $\eta_{2}(t)$ in (29) and (30), respectively. One can easily verify that
$\eta_{1}(t)=\left\{\begin{array}[]{cc}\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})&\textrm{as $t\to 0$}\\ \frac{4}{3\|A\|_{3}^{\frac{3}{2}}}t^{\frac{3}{2}}+O(t)&\textrm{as $t\to\infty$}\end{array}\right$. (70)
and
$\eta_{2}(t)=\left\{\begin{array}[]{cc}\frac{t}{4\|A\|}+O(t^{2})&\textrm{as $t\to 0$}\\ \frac{t}{\|A\|}+O(1)&\textrm{as $t\to\infty$}\end{array}\right.$. (73)
Figure 6: Inequality (31) in Theorem 2 is tighter than HWI for both sufficiently small and large values of $t$. For the symmetric matrix $A$ given in (51), Panel (a) and Panel (b) compare the two bounds for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We have also included the $m_{\infty}$-bound in Corollary 1 for comparison.
Therefore,
$\min\{\eta_{1}(t),\eta_{2}(t)\}=\left\{\begin{array}[]{cc}\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})&\textrm{as $t\to 0$}\\ \frac{t}{\|A\|}+O(1)&\textrm{as $t\to\infty$}\end{array}\right.$. (76)
This shows that $\min\{\eta_{1}(t),\eta_{2}(t)\}$ behaves similarly to the term $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$ for sufficiently small and large values of $t$. Since the coefficient $\kappa^{\prime}\approx 0.187$ that sits behind $\min\{\eta_{1}(t),\eta_{2}(t)\}$ is larger than the coefficient $\kappa\approx 0.145$ behind $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$ in HWI, it follows that (31) is tighter than HWI for those extreme values of $t$.
II-D Upper bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through $n$ and $\|A\|$
Let the matrix $A$ be positive-semidefinite. We present five upper bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through the dimension $n$ and the operator norm $\|A\|$. The first candidate is the $m_{\infty}$-bound in Corollary 1 which in fact holds for every symmetric matrix $A$. For the second and third candidates, we consider relaxed versions of HWI and LMI, respectively. Recall the inequality
$\|A\|_{2}\leq\sqrt{n}\|A\|.$ (77)
The exponents in HWI and LMI are increasing in terms of $\|A\|_{2}$. This is clear in case of HWI and its relaxed version in (32) immediately follows. As for LMI in (8), note that $\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2}$ is positive and decreasing in terms of $\|A\|_{2}$ and hence, $(\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2})^{2}$ becomes increasing in terms of $\|A\|_{2}$. This together with (77) gives
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\bigg(-\Big(\frac{\sqrt{n\|A\|^{2}+2\|A\|t}-\sqrt{n}\|A\|\,}{2\|A\|}\Big)^{2}\bigg).$ (78)
Factoring $\sqrt{n}\|A\|$ in $\sqrt{n\|A\|^{2}+2\|A\|t}-\sqrt{n}\|A\|$ results in the relaxed LMI in (33). As stated in Proposition 4, the $m_{\infty}$-bound is always tighter than both the relaxed HW and the relaxed LM inequalities. One can also consider the possibility of relaxing the augmented and optimal LM inequalities in (47) and (48), respectively. This requires that we replace the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$ in (40) by $\frac{t\|A\|}{(\sqrt{n}\|A\|)^{2}}=\frac{t}{n\|A\|}$. Numerical results indicate that the $m_{\infty}$-bound is still tighter that these relaxed inequalities and hence, we do not pursue this direction.
Next, we present a different upper bound on $\Pr(\boldsymbol{\Delta}>t)$ for an arbitrary symmetric matrix $A$ whose largest eigenvalue is nonzero. This will lead us to a fourth candidate that depends on the positive-semidefinite matrix $A$ only through $n$ and $\|A\|$. By a special case of the Min-Max Theorem in matrix analysis [9],
$\boldsymbol{x}^{T}A\boldsymbol{x}\leq\lambda_{\max}(A)\boldsymbol{x}^{T}\boldsymbol{x},$ (79)
where $\lambda_{\max}(A)$ is the largest eigenvalue of $A$. Note that $\boldsymbol{x}^{T}\boldsymbol{x}$ is a $\chi^{2}_{n}$ random variable. [^15] Chi-square with $n$ degrees of freedom. We consider two cases. If $\lambda_{\max}(A)>0$, then
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $\Pr(\lambda_{\max}(A)\boldsymbol{x}^{T}\boldsymbol{x}-\mathrm{Tr}(A)>t)$ (80)
$=$ $\Pr\Big(\boldsymbol{x}^{T}\boldsymbol{x}>\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)}\Big)$
$=$ $1-F_{\chi^{2}_{n}}\Big(\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)}\Big),$
where $F_{\chi^{2}_{n}(\cdot)}$ is the Cumulative Distribution Function (CDF) for $\chi^{2}_{n}$. If $\lambda_{\max}(A)<0$, then
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $\Pr(\lambda_{\max}(A)\boldsymbol{x}^{T}\boldsymbol{x}-\mathrm{Tr}(A)>t)$ (81)
$=$ $\Pr\Big(\boldsymbol{x}^{T}\boldsymbol{x}<\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)}\Big)$
$=$ $F_{\chi^{2}_{n}}\Big(\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)}\Big).$
In one line, we have
$\Pr(\boldsymbol{\Delta}>t)\leq\left\{\begin{array}[]{cc}1-F_{\chi^{2}_{n}}(\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)})&\lambda_{\max}(A)>0\\ F_{\chi^{2}_{n}}(\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)})&\lambda_{\max}(A)<0\end{array}\right.$. (84)
We will refer to the upper bound on the right side in (84) as the $\chi^{2}$-upper bound and to the whole inequality as the $\chi^{2}$-inequality. If $A$ is positive-semidefinite, then $\lambda_{\max}(A)=\|A\|$ and $\mathrm{Tr}(A)\geq\lambda_{\max}(A)$. Therefore, $\frac{\mathrm{Tr}(A)+t}{\lambda_{\max}(A)}\geq\frac{\|A\|+t}{\|A\|}=1+\frac{t}{\|A\|}$ and we derive the version of the $\chi^{2}$-bound in (34) as our fourth candidate. It is well-known that the CDF of $\chi^{2}_{n}$ is given by
$F_{\chi^{2}_{n}}(x)=\frac{\gamma(\frac{n}{2},\frac{x}{2})}{\Gamma(\frac{n}{2})},$ (85)
where $\gamma(a,z)=\int_{0}^{z}v^{a-1}e^{-v}dv$ is the lower incomplete gamma function and $\Gamma(a)=\int_{0}^{\infty}v^{a-1}e^{-v}dv$ is the gamma function. For simplicity and to gain some insight, let us assume that $n$ is an even integer during the rest of this discussion. In this case, $F_{\chi^{2}_{n}}(x)$ admits a closed-form expression given by [^16] One can derive (86) by using the formula $\int v^{k}e^{-v}dv=-\sum_{i=0}^{k}\frac{k!}{i!}v^{i}e^{-v}+C$ and the fact that $\Gamma(k)=(k-1)!$ for positive integer $k$.
$F_{\chi^{2}_{n}}(x)=1-\sum_{i=0}^{\frac{n}{2}-1}\frac{x^{i}e^{-\frac{x}{2}}}{2^{i}i!}.$ (86)
By (34) and (86),
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $\sum_{i=0}^{\frac{n}{2}-1}\frac{1}{2^{i}i!}\Big(1+\frac{t}{\|A\|}\Big)^{i}e^{-\frac{1}{2}(1+\frac{t}{\|A\|})}$ (87)
$=$ $\bigg(\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-1}\frac{1}{2^{i}i!}\Big(1+\frac{t}{\|A\|}\Big)^{i}\bigg)e^{-\frac{t}{2\|A\|}}.$
Both the $m_{\infty}$-bound in (52) and the $\chi^{2}$-bound in (87) depend on $A$ and $t$ through the dimension $n$ and the ratio
$r=\frac{t}{\|A\|}.$ (88)
We can write the $m_{\infty}$-inequality as
$\Pr(\boldsymbol{\Delta}>t)\leq P^{(1)}_{n}(r)e^{-\frac{r}{2}}$ (89)
and the $\chi^{2}$-inequality as
$\Pr(\boldsymbol{\Delta}>t)\leq P_{n}^{(2)}(r)e^{-\frac{r}{2}},$ (90)
where the polynomials $P_{n}^{(1)}(r)$ and $P^{(2)}_{n}(r)$ are defined by
$P_{n}^{(1)}(r)=\Big(1+\frac{r}{n}\Big)^{\frac{n}{2}}$ (91)
and
$P_{n}^{(2)}(r)=\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-1}\frac{(1+r)^{i}}{2^{i}i!},$ (92)
respectively. Any comparison between the two bounds boils down to comparing the two polynomials in (91) and (92). The first claim in Proposition 5 is that $P^{(2)}_{n}(r)<P_{n}^{(1)}(r)$ for every $r\geq 0$ when $n=2,4,6$. The second claim is that if $n$ is an even integer greater than or equal to $8$, then $P_{n}^{(1)}(r)<P_{n}^{(2)}(r)$ if and only if $r$ lies between the key numbers $r_{n},r^{\prime}_{n}$. These numbers are in fact the only two positive roots for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ as verified in the proof of Proposition 5 in Section IX.
Figure 7: Panels (a) and (b) present the plots for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for $n=8$ and $n=10$, respectively.
Fig. 7 in panels (a) and (b) presents graphs of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for $n=8$ and $n=10$, respectively. Approximate values for the numbers $r_{n},r^{\prime}_{n}$ are also provided on these plots. Fig. 8 compares the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$.
Figure 8: Comparison between the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$. The formulas for these four bounds are entirely functions of the dimension $n$ and the ratio $r$ in (88).
We close this section by recalling the well-known large deviations lower bound on the CDF for a $\chi^{2}_{n}$ random variable given by
$F_{\chi^{2}_{n}}(x)\geq 1-\mathrm{exp}\Big(-\frac{x}{2}+\frac{n}{2}\Big(1-\ln\frac{n}{x}\Big)\Big),\,\,\,\,\,x\geq n.$ (93)
This bound is a consequence of Markov’s inequality and a proof can be found in [10]. We emphasize that (93) only holds for $x\geq n$. Then one can loosen the $\chi^{2}$-bound in (34) as
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $1-F_{\chi^{2}_{n}}(1+r)$ (94)
$\leq$ $\mathrm{exp}\Big(-\frac{1+r}{2}+\frac{n}{2}\Big(1-\ln\frac{n}{1+r}\Big)\Big)$
$=$ $\frac{1}{\sqrt{e}}\Big(\frac{e(1+r)}{n}\Big)^{\frac{n}{2}}e^{-\frac{r}{2}},\,\,\,\,\,1+r\geq n.$
This is the large deviations bound introduced earlier in (35). It is always weaker than the $m_{\infty}$-bound for every $n\geq 2$. This is shown at the end of the proof for Proposition 4 in Section VIII.
III Proof of Proposition 1
The indispensable starting point in proving HWI is the exponential Markov inequality which gives
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-st}\mathbb{E}[e^{s\boldsymbol{\Delta}}],$ (95)
where $s\geq 0$ is arbitrary. Let $A=UDU^{T}$ be the eigenvalue decomposition of $A$ where $U$ is an orthogonal matrix and the diagonal matrix $D$ carries the (real) eigenvalues $\lambda_{1},\cdots,\lambda_{n}$ of $A$. Then
$\boldsymbol{\Delta}=\sum_{i=1}^{n}\lambda_{i}(\boldsymbol{\xi}^{2}_{i}-1),$ (96)
where $\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{n}$ are independent standard normal random variables that are the entries of the random vector $\underline{\boldsymbol{\xi}}=U^{T}\underline{\boldsymbol{x}}$. By independence,
$\mathbb{E}[e^{s\boldsymbol{\Delta}}]=\prod_{i=1}^{n}e^{-s\lambda_{i}}\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}].$ (97)
Since $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $2s\lambda_{i}<1$, we get
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{-\frac{1}{2}\ln(1-2s\lambda_{i})},$ (98)
for every $s\geq 0$ such that $2s\lambda_{i}<1$ for all $1\leq i\leq n$. Let $a>0$ and $0<b<1$ be such that (10) holds. The bound in (98) can be loosened as
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+a(2s\lambda_{i})^{2})}$ (99)
$=$ $\exp\Big(-st+2as^{2}\sum_{i=1}^{n}\lambda_{i}^{2}\Big)$
$=$ $\exp\big(-st+2a\|A\|_{2}^{2}s^{2}\big),$
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The minimum value for the function
$f(s)=-st+2a\|A\|_{2}^{2}s^{2}$
over $0\leq s\leq\frac{b}{2\|A\|}$ occurs at $s_{opt}=\min\{s_{0},\frac{b}{2\|A\|}\}$ where $s_{0}=\frac{t}{4a\|A\|_{2}^{2}}$. We consider two cases.
  1. 1.
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then
    $f(s_{opt})$ $=$ $f(s_{0})$ (100)
    $=$ $-\frac{t^{2}}{4a\|A\|_{2}^{2}}+2a\|A\|_{2}^{2}\Big(\frac{t}{4a\|A\|_{2}^{2}}\Big)^{2}$
    $=$ $-\frac{t^{2}}{8a\|A\|_{2}^{2}}.$
  2. 2.
    Let $s_{0}>\frac{b}{2\|A\|}$. Then
    $f(s_{opt})$ $=$ $f\Big(\frac{b}{2\|A\|}\Big)$ (101)
    $=$ $-\frac{bt}{2\|A\|}+2a\|A\|_{2}^{2}\Big(\frac{b}{2\|A\|}\Big)^{2}$
    $\stackrel{{\scriptstyle}}{{<}}$ $-\frac{bt}{2\|A\|}+2a\|A\|_{2}^{2}\times s_{0}\times\frac{b}{2\|A\|}$
    $=$ $-\frac{bt}{2\|A\|}+\frac{t}{2}\times\frac{b}{2\|A\|}$
    $=$ $-\frac{bt}{4\|A\|}.$
By (100) and (101),
$f(s_{opt})$ $\leq$ $\max\Big\{-\frac{t^{2}}{8a\|A\|_{2}^{2}},-\frac{bt}{4\|A\|}\Big\}$ (102)
$=$ $-\min\Big\{\frac{t^{2}}{8a\|A\|_{2}^{2}},\frac{bt}{4\|A\|}\Big\}$
$\leq$ $-\kappa(a,b)\min\Big\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\Big\},$
where
$\kappa(a,b)=\min\Big\{\frac{1}{8a},\frac{b}{4}\Big\}.$ (103)
Next, let us explore the constants $a,b>0$. Reference [4] considers $a=1$ and $b=\frac{1}{2}$. This results in $\kappa=\kappa(1,0.5)=0.125$. But, one can do better. We need the special case of the following lemma for $m=1$.
Lemma 1.
Let $m\geq 1$ be an integer. For every $0<b<1$, the smallest $a$ (tightest upper bound) for which (23) holds is $a=\theta_{m}(b)$ as defined in (11). Equivalently, for every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (10) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (10) holds.
Proof.
See Appendix A. ∎
Applying Lemma 1 with $m=1$ results in
$\kappa=\kappa(\theta_{1}(b),b)=\min\Big\{\frac{1}{8\theta_{1}(b)},\frac{b}{4}\Big\}.$ (104)
for arbitrary $0<b<1$. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $2b\theta_{1}(b)=1$. This equation is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 9.
Figure 9: The maximum value for $\kappa(\theta_{1}(b),b)=\min\big\{\frac{1}{8\theta_{1}(b)},\frac{b}{4}\big\}$ is achieved where $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$. This is solved for $b\approx 0.583$ and the attained maximum value is approximately $0.1457$.
IV Proof of Proposition 2
Since $A$ is positive-semidefinite, its eigenvalues are all nonnegative. We begin with the inequality in (98). Assume (14) holds for some $a,b>0$ such that $ab<1$. Then
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+\frac{(2s\lambda_{i})^{2}}{2(1-2as\lambda_{i})})}$ (105)
$=$ $e^{-st}\prod_{i=1}^{n}e^{\frac{s^{2}\lambda^{2}_{i}}{1-2as\lambda_{i}}}$
$=$ $\exp\Big(-st+\sum_{i=1}^{n}\frac{s^{2}\lambda_{i}^{2}}{1-2as\lambda_{i}}\Big),$
for every $s\geq 0$ such that $0\leq 2s\lambda_{i}\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. Since $ab<1$ by assumption, then
$0<1-ab\leq 1-2as\lambda_{i},$ (106)
for every $1\leq i\leq n$. Therefore, $0<1-2as\|A\|\leq 1-2as\lambda_{i}$ for every $1\leq i\leq n$ and we can loosen the upper bound in (105) as
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $\exp\Big(-st+\sum_{i=1}^{n}\frac{s^{2}\lambda_{i}^{2}}{1-2as\|A\|}\Big)$ (107)
$=$ $\exp\Big(-st+\frac{\|A\|_{2}^{2}s^{2}}{1-2as\|A\|}\Big),\,\,\,\,\,0\leq s\leq\frac{b}{2\|A\|}.$
Let $\alpha$ and $\beta$ be as in (15) and call the exponent on the right side of (107) by $f(s)$, i.e.,
$f(s)=-st+\frac{\beta s^{2}}{1-2a\alpha s}.$ (108)
It is straightforward to check that the function $f$ is convex (concave upward) over the entire interval $(-\infty,\frac{1}{2a\alpha})$ on the left side of its vertical asymptote $s=\frac{1}{2a\alpha}$. Moreover, it achieves an absolute minimum value over that interval at
$s_{0}=\frac{1}{2a\alpha}\Bigg(1-\frac{1}{\big(1+\frac{2a\alpha t}{\beta}\big)^{\frac{1}{2}}}\Bigg).$ (109)
Note that $\frac{b}{2\alpha}<\frac{1}{2a\alpha}$ due to the assumption $ab<1$. It follows that the expression on the right side of (107) achieves its minimum value over $0\leq s\leq\frac{b}{2\alpha}$ at
$s_{opt}=\min\Big\{s_{0},\frac{b}{2\alpha}\Big\}.$ (110)
We consider two cases.
  1. 1.
    Let $s_{0}\leq\frac{b}{2\alpha}$, or equivalently,
    $\frac{\alpha t}{\beta}\leq\frac{1}{2a}\Big(\frac{1}{(1-ab)^{2}}-1\Big).$ (111)
    Then
    $f(s_{opt})=f(s_{0})=-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\bigg(\Big(1+\frac{2a\alpha t}{\beta}\Big)^{\frac{1}{2}}-1\bigg),$ (112)
    where we omit the simple algebra.
  2. 2.
    Let $s_{0}>\frac{b}{2\alpha}$. Then
    $f(s_{opt})=f\Big(\frac{b}{2\alpha}\Big)=-\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}$ (113)
Let us denote the expression on the right side of (111) by $c$, i.e.,
$c=\frac{1}{2a}\Big(\frac{1}{(1-ab)^{2}}-1\Big).$ (114)
We have shown that
$\Pr(\boldsymbol{\Delta}>t)\leq e^{f(s_{opt})}$ (115)
and
$f(s_{opt})=\left\{\begin{array}[]{cc}-\frac{t}{2a\alpha}+\frac{\beta}{2a^{2}\alpha^{2}}\Big(\big(1+\frac{2a\alpha t}{\beta}\big)^{\frac{1}{2}}-1\Big)&\frac{\alpha t}{\beta}\leq c\\ -\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}&\frac{\alpha t}{\beta}>c\end{array}\right.$. (118)
The next lemma explores possibilities for $a,b$. It addresses an inequality for which (14) is a special case.
Lemma 2.
Let $m\geq 1$ be an integer. For every $\frac{m+1}{m+2}<a<1$, the inequality
$-\ln(1-x)\leq x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+\frac{x^{m+1}}{(m+1)(1-ax)},\,\,\,\,\,0\leq x\leq b,$ (119)
holds with
$b=\frac{m+1-(m+2)a}{a(m-(m+1)a)}.$ (120)
Proof.
See Appendix B. ∎
Applying Lemma 2 with $m=1$, we obtain
$b=\frac{3a-2}{a(2a-1)},\,\,\,\,\,\frac{2}{3}<a<1.$ (121)
Note that $ab=\frac{3a-2}{2a-1}<1$ as required in (106). Moreover, $1-ab=\frac{1-a}{2a-1}$ and the threshold $c$ in (114) is given by
$c=\frac{3a-2}{2(1-a)^{2}}.$ (122)
V Proof of Proposition 3
To see how (17) implies HWI, let us write HWI in (2) as
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}},$ (123)
where $\alpha,\beta$ are given in (15). Note
$\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}=\left\{\begin{array}[]{cc}\frac{t^{2}}{\beta}&\frac{\alpha t}{\beta}\leq 1\\ \frac{t}{\alpha}&\frac{\alpha t}{\beta}>1\end{array}\right.$. (126)
We aim to show that there exists $\kappa>0$ such that the upper bound in (17) is less than or equal to the upper bound in (123), or equivalently, $\Lambda(t,a)\leq-\kappa\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}$. We want this to hold regardless of $t,\alpha,\beta$ and $\kappa$ to be as large as possible. Thus, we are looking to compute
$\kappa=\inf_{t,\alpha,\beta>0}-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (127)
Interestingly, $\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio $\rho=\frac{\alpha t}{\beta}$ in (40). For convenience in comparing the right sides in (20) and (126), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible interval $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
$a_{0}=\frac{7-\sqrt{17}}{4}\approx 0.7192.$ (128)
One can easily check that the corresponding value for $b$ is $b_{0}=\frac{3a_{0}-2}{a_{0}(2a_{0}-1)}=\frac{1}{2}$. We have
$-\frac{\Lambda(t,a_{0})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a_{0}\rho}-\frac{1}{2a_{0}^{2}\rho^{2}}\big(\sqrt{1+2a_{0}\rho}-1\big)&\rho\leq 1\\ \frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})\rho}&\rho>1\end{array}\right.$. (131)
Fig. 10 shows the graph of the function of $\rho$ that sits on the right side of (131). It achieves its absolute minimum value at $\rho=1$. It follows that the infimum in (127) is given by
$\frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})}=\frac{9-\sqrt{17}}{32}\approx 0.152.$ (132)
Figure 10: Graph of the function of $\rho$ given on the right side of (131). It achieves its absolute minimum value at $\rho=1$.
It is shown in Appendix C that the choice of $a=a_{0}$ in (128) indeed results in the largest value for $\kappa$ and that it was not just a choice of convenience.
Finally, we address LMI itself. To see why (7) and (8) are equivalent, call $t^{\prime}=2\sqrt{\beta t}+2\alpha t$. Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LMI in (7) becomes
$\Pr(\boldsymbol{\Delta}>t)\leq\exp\left(-\Big(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha}\Big)^{2}\right).$ (133)
This inequality implies HWI in (123) with
$\kappa=\inf_{t,\alpha,\beta>0}-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (134)
Once again, the ratio on the right side of (134) depends entirely on $\rho=\frac{\alpha t}{\beta}$. It is given by
$-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}(\frac{\sqrt{1+2\rho}-1}{2\rho})^{2}&\rho\leq 1\\ \frac{(\sqrt{1+2\rho}-1)^{2}}{4\rho}&\rho>1\end{array}\right.$. (137)
This function of $\rho$ achieves it absolute minimum value of $1-\frac{\sqrt{3}}{2}$ at $\rho=1$.
VI Proof of Theorem 1
We begin with the bound in (98). By Lemma 1 in Section III, for every integer $m\geq 1$ and $0<b<1$, the inequality in (23) is the tightest for $a=\theta_{m}(b)$. Then
$\Pr(\boldsymbol{\Delta}>t)$ $\leq$ $e^{-st}\prod_{i=1}^{n}e^{-s\lambda_{i}}e^{\frac{1}{2}(2s\lambda_{i}+a|2s\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{(2s\lambda_{i})^{k}}{k})}$ (138)
$=$ $\exp\Big(-st+a2^{m}s^{m+1}\sum_{i=1}^{n}|\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{2^{k-1}s^{k}}{k}\sum_{i=1}^{n}\lambda_{i}^{k}\Big)$
$=$ $\exp\Big(-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}+\sum_{k=2}^{m}\frac{2^{k-1}s^{k}\mathrm{Tr}(A^{k})}{k}\Big),$
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (138) is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
$s_{0}=\frac{1}{2}\left(\frac{t}{(m+1)a\|A\|_{m+1}^{m+1}}\right)^{\frac{1}{m}}.$ (139)
We compute the right side of (138) at $s=\hat{s}_{opt}$ defined by
$\hat{s}_{opt}=\min\Big\{s_{0},\frac{b}{2\|A\|}\Big\}.$ (140)
Then the third term in the exponent in (138) is bounded by
$\sum_{k=2}^{m}\frac{2^{k-1}\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})}{k}$ $\leq$ $\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}|\mathrm{Tr}(A^{k})|}{k}$ (141)
$=$ $\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{|\mathrm{Tr}(A^{k})|}{\|A\|^{k}}$
$\leq$ $\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$
where the first step is due to $\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})\leq|\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})|$ and $|\hat{s}_{opt}|=\hat{s}_{opt}\leq\frac{b}{2\|A\|}$ and the last step is due to [^17] For a symmetric matrix $A$, the operator norm $\|A\|$ is the maximum value among the absolute values of eigenvalues of $A$.
$|\mathrm{Tr}(A^{k})|=\big|\sum_{i=1}^{n}\lambda_{i}^{k}\big|\leq\sum_{i=1}^{n}|\lambda_{i}|^{k}\leq n(\max_{1\leq i\leq n}|\lambda_{i}|)^{k}=n\|A\|^{k}.$ (142)
Next, we bound the sum of the first two terms in the exponent, i.e., $\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases.
  1. 1.
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then $\hat{s}_{opt}=s_{0}$ and simple algebra shows that
    $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}=-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},$ (143)
    where $\kappa_{m,1}$ is defined by
    $\kappa_{m,1}=\frac{1-\frac{1}{m+1}}{2(m+1)^{\frac{1}{m}}a^{\frac{1}{m}}}.$ (144)
  2. 2.
    Let $s_{0}>\frac{b}{2\|A\|}$. Then $\hat{s}_{opt}=\frac{b}{2\|A\|}$ and we have
    $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$ $=$ $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}\times\hat{s}_{opt}^{m}\|A\|_{m+1}^{m+1}$ (145)
    $\leq$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times s_{0}^{m}\|A\|_{m+1}^{m+1}$
    $=$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times\frac{t}{2^{m}(m+1)a\|A\|_{m+1}^{m+1}}\|A\|_{m+1}^{m+1}$
    $=$ $-\kappa_{m,2}\frac{t}{\|A\|},$
    where the second step is due to $\hat{s}_{opt}\leq s_{0}$ and $\kappa_{m,2}$ is defined by
    $\kappa_{m,2}=\frac{b}{2}\Big(1-\frac{1}{m+1}\Big).$ (146)
By (143) and (145),
$-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$ $\leq$ $\max\Big\{-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},-\kappa_{m,2}\frac{t}{\|A\|}\Big\}$ (147)
$=$ $-\min\Big\{\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\kappa_{m,2}\frac{t}{\|A\|}\Big\}$
$\leq$ $-\min\{\kappa_{m,1},\kappa_{m,2}\}\min\bigg\{\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg\}.$
The proof of Theorem 1 is now complete.
VII Proof of Theorem 2
Consider the bound in (138) for $m=2$, i.e.,
$\Pr(\boldsymbol{\Delta}>t)\leq e^{f(s)},$ (148)
where $f(s)$ is given by
$f(s)=-st+\beta s^{2}+4a\gamma s^{3},\,\,\,\,\,0\leq s\leq\frac{b}{2\alpha},$ (149)
$\alpha,\beta$ and $\gamma$ are defined in (28) and $a=\theta_{2}(b)$ for $0<b<1$. The function $f(s)$ achieves its global minimum value over $s\geq 0$ at $s=s_{0}$ given by
$s_{0}=\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{12a\gamma}.$ (150)
It follows that the minimum value for $f(s)$ over the interval $0\leq s\leq\frac{b}{2\alpha}$ is achieved at
$s_{opt}=\min\Big\{s_{0},\frac{b}{2\alpha}\Big\}.$ (151)
Two cases are considered.
  1. 1.
    Let $s_{0}\leq\frac{b}{2\alpha}$. Then $s_{opt}=s_{0}$ and $f(s_{opt})=-s_{0}t+\beta s_{0}^{2}+4a\gamma s_{0}^{3}$. We need the following lemma.
    Lemma 3.
    Let $u,v,w,w^{\prime}$ be positive numbers such that $w\geq w^{\prime}$ and $x=\frac{\sqrt{v^{2}+3uw}-v}{3w}$. Then
    $\hskip 28.45274pt-ux+vx^{2}+wx^{3}\leq\frac{1}{27w^{\frac{1}{2}}w^{\prime\frac{3}{2}}}(\sqrt{v^{2}+3uw^{\prime}}-v)(v\sqrt{v^{2}+3uw^{\prime}}-v^{2}-6uw^{\prime}).$ (152)
    Proof.
    See Appendix D. ∎
    Applying Lemma 3 with $x=s_{0}$, $u=t$, $v=\beta$ and $w=4a\gamma$ and noting that $w\geq w^{\prime}=\frac{4}{3}\gamma$ due to $a=\theta_{2}(b)\geq\frac{1}{3}$, we obtain
    $f(s_{opt})\leq\frac{\sqrt{3}}{144a^{\frac{1}{2}}\gamma^{2}}(\sqrt{\beta^{2}+4\gamma t}-\beta)(\beta\sqrt{\beta^{2}+4\gamma t}-\beta^{2}-8\gamma t).$ (153)
    Recalling the definition of $\eta_{1}(t)$ in (29), we can write (153) as
    $f(s_{opt})\leq-\frac{\sqrt{3}}{12a^{\frac{1}{2}}}\eta_{1}(t).$ (154)
  2. 2.
    Let $s_{0}>\frac{b}{2\alpha}$. Then $s_{opt}=\frac{b}{2\alpha}$ and
    $f(s_{opt})$ $=$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma s_{opt}^{3}$ (155)
    $=$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma s_{opt}\times s_{opt}^{2}$
    $\leq$ $-\frac{bt}{2\alpha}+\beta s_{opt}^{2}+4a\gamma\frac{b}{2\alpha}\times s_{0}^{2},$
    where the last step is due to $s_{opt}=\frac{b}{2\alpha}<s_{0}$. Next, we bound the second and third terms on the right side of (155). We write the second term as
    $\beta s_{opt}^{2}$ $=$ $\beta s_{opt}\times s_{opt}$ (156)
    $\leq$ $\beta\frac{b}{2\alpha}\times s_{0}$
    $=$ $\frac{\beta b}{2\alpha}\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{12a\gamma}$
    $\leq$ $\frac{\beta b}{24\alpha\gamma}\frac{\sqrt{\beta^{2}+12\times\frac{1}{3}\times\gamma t}-\beta}{\frac{1}{3}}$
    $=$ $\frac{\beta b}{8\alpha\gamma}(\sqrt{\beta^{2}+4\gamma t}-\beta),$
    where the penultimate step uses the fact that $a\mapsto\frac{\sqrt{\beta^{2}+12a\gamma t}-\beta}{a}$ is a decreasing function of $a$ and $a=\theta_{2}(b)\geq\frac{1}{3}$. Alternatively, we can bound $\beta s_{opt}^{2}$ as
    $\beta s_{opt}^{2}=\frac{\beta b^{2}}{4\alpha^{2}}\leq\frac{\beta b}{4\alpha^{2}},$ (157)
    due to $b^{2}\leq b$ as $0<b<1$. By (156) and (157), we get
    $\beta s_{opt}^{2}\leq\frac{\beta b}{4\alpha}\min\Big\{\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t}-\beta}{2\gamma}\Big\}.$ (158)
    Regarding the third term on the right side of (155), note that $s_{0}\leq\frac{\sqrt{\beta^{2}}+\sqrt{12a\gamma t}-\beta}{12a\gamma}=\sqrt{\frac{t}{12a\gamma}}$ due to the trivial inequality $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for $x,y\geq 0$. Then
    $4a\gamma\frac{b}{2\alpha}\times s_{0}^{2}\leq 4a\gamma\frac{b}{2\alpha}\times\frac{t}{12a\gamma}=\frac{bt}{6\alpha}.$ (159)
    Putting (155), (158) and (159) together, we find
    $f(s_{opt})$ $\leq$ $-\frac{bt}{2\alpha}+\frac{\beta b}{4\alpha}\min\Big\{\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t}-\beta}{2\gamma}\Big\}+\frac{bt}{6\alpha}$ (160)
    $=$ $\frac{\beta b}{4\alpha}\min\Big\{\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t}-\beta}{2\gamma}\Big\}-\frac{bt}{3\alpha}$
    $=$ $-\frac{b}{3}\Big(\frac{t}{\alpha}-\frac{3\beta}{4\alpha}\min\Big\{\frac{1}{\alpha},\frac{\sqrt{\beta^{2}+4\gamma t}-\beta}{2\gamma}\Big\}\Big)$
    $=$ $-\frac{b}{3}\eta_{2}(t),$
    where $\eta_{2}(t)$ is defined in (30).
By (154) and (160),
$f(s_{opt})$ $\leq$ $\max\Big\{-\frac{\sqrt{3}}{12a^{\frac{1}{2}}}\eta_{1}(t),-\frac{b}{3}\eta_{2}(t)\Big\}$ (161)
$=$ $-\min\Big\{\frac{\sqrt{3}}{12a^{\frac{1}{2}}}\eta_{1}(t),\frac{b}{3}\eta_{2}(t)\Big\}$
$\leq$ $-\min\Big\{\frac{\sqrt{3}}{12a^{\frac{1}{2}}},\frac{b}{3}\Big\}\min\{\eta_{1}(t),\eta_{2}(t)\},$
where the last step is due to the fact that $\eta_{1}(t),\eta_{2}(t)\geq 0$. Finally, we select $0<b<1$ such that $\min\big\{\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}},\frac{b}{3}\big\}$ is maximized. This happens when $\frac{\sqrt{3}}{12(\theta_{2}(b))^{\frac{1}{2}}}=\frac{b}{3}$ or equivalently, $\frac{4}{\sqrt{3}}b(\theta_{2}(b))^{\frac{1}{2}}=1$ as promised.
VIII Proof of Proposition 4
Let us write the $m_{\infty}$-bound in (27) and the relaxed HW bound in (32) in terms of $r=\frac{t}{\|A\|}$. The $m_{\infty}$-bound becomes
$\Pr(\boldsymbol{\Delta}>t)\leq\Big(1+\frac{r}{n}\Big)^{\frac{n}{2}}e^{-\frac{r}{2}}=e^{-(\frac{r}{2}-\frac{n}{2}\ln(1+\frac{r}{n}))}.$ (162)
and the relaxed HWI becomes
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\{\frac{r^{2}}{n},r\}},\,\,\,\,\,\kappa=\frac{9-\sqrt{17}}{32}.$ (163)
The $m_{\infty}$-bound is tighter than the relaxed HWI if and only if
$\frac{r}{2}-\frac{n}{2}\ln\Big(1+\frac{r}{n}\Big)>\kappa\min\Big\{\frac{r^{2}}{n},r\Big\},\,\,\,\,\,r>0.$ (164)
We consider two cases.
  1. 1.
    Let $0<r\leq n$. Then $\min\{\frac{r^{2}}{n},r\}=\frac{r^{2}}{n}$ and (164) becomes
    $\frac{r}{2}-\frac{n}{2}\ln\Big(1+\frac{r}{n}\Big)>\frac{\kappa r^{2}}{n}.$ (165)
    Dividing both sides by $\frac{n}{2}$ and rearranging terms,
    $\ln\Big(1+\frac{r}{n}\Big)<\frac{r}{n}\Big(1-\frac{2\kappa r}{n}\Big).$ (166)
    Denoting $x=\frac{r}{n}$, we need to show that $\ln(1+x)<x(1-2\kappa x)$ for all $0<x\leq 1$. Define $f(x)=x(1-2\kappa x)-\ln(1+x)$. Then $f^{\prime}(x)=1-4\kappa x-\frac{1}{1+x}=\frac{(1-4\kappa)x-4\kappa x^{2}}{1+x}$. The numerator is the quadratic function $q(x)=(1-4\kappa)x-4\kappa x^{2}$. Solving $q(x)=0$, we see that $f(x)$ has only one critical number inside the interval $(0,1)$ given by $x_{0}=\frac{1-4\kappa}{4\kappa}=\frac{1+\sqrt{17}}{8}\approx 0.6404$. Moreover, $f^{\prime}(x)>0$ for $0<x<x_{0}$ and $f^{\prime}(x)<0$ for $x_{0}<x<1$. By the first derivative test, the function $f$ increases over $(0,x_{0})$, achieves a local maximum value at $x=x_{0}$ and then decreases over $(x_{0},1)$. As such, $f$ achieves it absolute minimum value over the interval $[0,1]$ either at $x=0$ or at $x=1$. But, $f(0)=0$ and $f(1)=1-2\kappa-\ln 2\approx 0.002>0$. It follows that $f(x)>0$ over $(0,1]$ as desired.
  2. 2.
    Let $r>n$. Then $\min\{\frac{r^{2}}{n},r\}=r$ and (164) becomes $\frac{r}{2}-\frac{n}{2}\ln(1+\frac{r}{n})>\kappa r$. Dividing both sides by $\frac{n}{2}$ and rearranging terms, we get the inequality
    $\ln\Big(1+\frac{r}{n}\Big)<(1-2\kappa)\frac{r}{n}.$ (167)
    Denoting $x=\frac{r}{n}$, we show that $\ln(1+x)<(1-2\kappa)x$ for every $x>1$. Define the function $f(x)=(1-2\kappa)x-\ln(1+x)$. Then $f^{\prime}(x)=1-2\kappa-\frac{1}{1+x}>1-2\kappa-\frac{1}{2}=\frac{\sqrt{17}-1}{16}>0$ where we used $\frac{1}{1+x}>-\frac{1}{2}$ for $x>1$. Since $f(1)=1-2\kappa-\ln 2\approx 0.002>0$ and $f$ is increasing over $1,\infty)$, it follows that $f(x)>0$ for every $x>1$ as desired.
Next, we look at the relaxed LMI in (33). Writing the bound in terms of $r=\frac{t}{\|A\|}$ and expanding the complete square in the exponent, we have
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\frac{r}{2}+\frac{n}{2}(\sqrt{1+\frac{2r}{n}}-1)}.$ (168)
Therefore, the upper bound in the $m_{\infty}$-inequality in (162) is less than the upper bound in the relaxed LMI if and only if
$\ln\Big(1+\frac{r}{n}\Big)<\sqrt{1+\frac{2r}{n}}-1.$ (169)
But, this is true for all $r>0$ thanks to the inequality $\ln(1+x)<\sqrt{1+2x}-1$ for every $x>0$. To see this, let $f(x)=\sqrt{1+2x}-1-\ln(1+x)$. Then $f^{\prime}(x)=\frac{1}{\sqrt{1+2x}}-\frac{1}{1+x}$. But, $\sqrt{1+2x}<1+x$ for every $x>0$. As such, $f^{\prime}(x)>0$ for every $x>0$. Since $f(0)=0$ and $f$ is increasing over $0,\infty)$, then $f(x)>f(0)=0$ for every $x>0$.
Finally, we verify that the $m_{\infty}$-bound always beats the large deviations bound. Call $x=\frac{1+r}{n}\geq 1$. The bound in (94) is looser than that in (89) if and only if
$\Big(1+\frac{r}{n}\Big)^{\frac{n}{2}}<\frac{1}{\sqrt{e}}\Big(\frac{e(1+r)}{n}\Big)^{\frac{n}{2}}$ (170)
$\iff$ $\Big(1+\frac{nx-1}{n}\Big)^{\frac{n}{2}}<\frac{1}{\sqrt{e}}(ex)^{\frac{n}{2}}$
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ $1+\frac{nx-1}{n}<\frac{1}{e^{\frac{1}{n}}}ex$
$\iff$ $1-\frac{1}{n}+x<e^{1-\frac{1}{n}}x$
$\iff$ $1-\frac{1}{n}<(e^{1-\frac{1}{n}}-1)x,$
where in ()(*) we raised both sides to the power of $\frac{2}{n}$. Using the inequality $z<e^{z}-1$ for $z\neq 0$, we have $1-\frac{1}{n}<e^{1-\frac{1}{n}}-1$ for every $n\geq 2$. Therefore, (170) is certainly true due to $x\geq 1$.
IX Proof of Proposition 5
Recall the polynomials $P_{n}^{(1)}(r)$ and $P_{n}^{(2)}(r)$ in (91) and (92), respectively. It is straightforward to check that $P^{(2)}_{n}(r)<P_{n}^{(1)}(r)$ for all $r\geq 0$ if $n=2,4,6$. In fact,
$P_{2}^{(1)}(r)-P_{2}^{(2)}(r)=\frac{r}{2}+1-\frac{1}{\sqrt{e}},$ (171)
$P_{4}^{(1)}(r)-P_{4}^{(2)}(r)=\frac{r^{2}}{16}+\Big(\frac{1}{2}-\frac{1}{2\sqrt{e}}\Big)r+1-\frac{3}{2\sqrt{e}}$ (172)
and
$P_{6}^{(1)}(r)-P_{6}^{(2)}(r)=\frac{r^{3}}{216}+\Big(\frac{1}{12}-\frac{1}{8\sqrt{e}}\Big)r^{2}+\Big(\frac{1}{2}-\frac{3}{4\sqrt{e}}\Big)r+1-\frac{13}{8\sqrt{e}}.$ (173)
The polynomials on the right sides in (171), (172) and (173) all have positive coefficients. This proves supremacy of the $\chi^{2}$-bound over the $m_{\infty}$-bound when $n=2,4,6$. If $n$ is an even integer greater than or equal to $8$, the proof strategy is to first show that the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ has at least two positive roots and then proceed to prove that $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ can not have more than two positive roots. We observe that $P_{n}^{(1)}(0)>P_{n}^{(2)}(0)$ regardless of the value of $n$. In fact,
$P_{n}^{(1)}(0)=1,\,\,\,\,\,P_{n}^{(2)}(0)=\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-1}\frac{1}{2^{i}i!}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\infty}\frac{(\frac{1}{2})^{i}}{i!}=1.$ (174)
Since $\deg(P_{n}^{(1)})=\frac{n}{2}$ is larger than $\deg(P_{n}^{(2)})=\frac{n}{2}-1$, we also have $P_{n}^{(1)}(r)>P_{n}^{(2)}(r)$ for all sufficiently large $r$. If we can find a positive number $r^{*}$ such that $P_{n}^{(1)}(r^{*})<P_{n}^{(2)}(r^{*})$, then the Intermediate Value Theorem (IVT) implies that there must exist numbers $0<r_{n}<r^{*}<r^{\prime}_{n}$ such that $P_{n}^{(1)}(r_{n})=P_{n}^{(2)}(r_{n})$ and $P_{n}^{(1)}(r^{\prime}_{n})=P_{n}^{(2)}(r^{\prime}_{n})$, i.e., $r_{n},r^{\prime}_{n}$ are zeros of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. We verify that the choice $r^{*}=1$ works for every even integer $n\geq 8$. Let us state this as a lemma.
Lemma 4.
Let the polynomials $P_{n}^{(1)}(r)$ and $P_{n}^{(2)}(r)$ be as in (91) and (92), respectively. Then $P_{n}^{(1)}(1)<P_{n}^{(2)}(1)$ for every even integer $n\geq 8$, i.e.,
$\Big(1+\frac{1}{n}\Big)^{\frac{n}{2}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-1}\frac{1}{i!},\,\,\,\,\,n\geq 8.$ (175)
Proof.
Denote $m=\frac{n}{2}$. We need to show that
$\Big(1+\frac{1}{2m}\Big)^{m}<\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{i!},\,\,\,\,\,m\geq 4.$ (176)
The cases $4\leq m\leq 49$ and $m\geq 50$ are addressed separately. [^18] The reason becomes clear by the end of proof for Lemma 4. Fig. 11 depicts the plots for $(1+\frac{1}{2m})^{m}$ and $\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{i!}$ in terms of integers $4\leq m\leq 49$. We see that (176) holds for $4\leq m\leq 49$.
Figure 11: Plots for $(1+\frac{1}{2m})^{m}$ and $\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{i!}$ in terms of $4\leq m\leq 49$.
During the rest of proof, we assume $m\geq 50$. Let $\boldsymbol{x}\sim\mathrm{Poisson}(1)$ be a Poisson random variable with parameter $1$. Then
$\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{i!}$ $=$ $\sqrt{e}\sum_{i=0}^{m-1}\frac{e^{-1}1^{i}}{i!}$ (177)
$=$ $\sqrt{e}\Pr(\boldsymbol{x}\leq m-1)$
$=$ $\sqrt{e}\big(1-\Pr(\boldsymbol{x}\geq m)\big)$
$=$ $\sqrt{e}\big(1-\Pr(\boldsymbol{x}^{2}\geq m^{2})\big)$
$\geq$ $\sqrt{e}\Big(1-\frac{\mathbb{E}[\boldsymbol{x}^{2}]}{m^{2}}\Big)$
$=$ $\sqrt{e}\Big(1-\frac{2}{m^{2}}\Big),$
where the penultimate step is due to Markov’s inequality and the last step is due to $\mathbb{E}[\boldsymbol{x}^{2}]=\mathrm{Var}(\boldsymbol{x})+(\mathbb{E}[\boldsymbol{x}])^{2}=1+1^{2}=2$. By (175) and (177), it suffices to show that
$\Big(1+\frac{1}{2m}\Big)^{m}<\sqrt{e}\Big(1-\frac{2}{m^{2}}\Big),\,\,\,\,\,m\geq 50.$ (178)
Define
$f(x)=\frac{1}{2}+\ln\Big(1-\frac{2}{x^{2}}\Big)-x\ln\Big(1+\frac{1}{2x}\Big),\,\,\,\,\,x\geq 50.$ (179)
The inequality in (178) can be written as $e^{f(m)}>1$ for $m\geq 50$. Thus, we aim to show that
$f(x)>0,\,\,\,\,\,x\geq 50.$ (180)
To achieve this goal, we prove that
$\lim_{x\to\infty}f(x)=0$ (181)
and
$f^{\prime}(x)<0,\,\,\,\,\,x\geq 50.$ (182)
One can easily check that (181) and (182) imply (180). Verifying (181) is an easy task. To prove (182), we use a similar method. We show that
$\lim_{x\to\infty}f^{\prime}(x)=0$ (183)
and
$f^{\prime\prime}(x)>0,\,\,\,\,\,x\geq 50.$ (184)
We have
$f^{\prime}(x)=\frac{4}{x(x^{2}-2)}+\frac{1}{2x+1}-\ln\Big(1+\frac{1}{2x}\Big).$ (185)
Then (183) clearly holds. To verify (184), we compute
$f^{\prime\prime}(x)=\frac{x^{5}-48x^{4}-52x^{3}+20x^{2}+36x+8}{x^{2}(x^{2}-2)^{2}(2x+1)^{2}},$ (186)
where we omit the tedious algebra. The denominator in (186) is clearly positive for every $x\geq 50$. Let us denote the numerator by $p(x)$, i.e.,
$p(x)=x^{5}-48x^{4}-52x^{3}+20x^{2}+36x+8.$ (187)
To show that $p(x)>0$ for every $x\geq 50$, let us look at the shifted polynomial $p(x+50)$. It is given by
$p(x+50)=x^{5}+202x^{4}+15348x^{3}+522220x^{2}+6862036x+6051808.$ (188)
Since all coefficients are positive in $p(x+50)$, we conclude that $p(x+50)\geq 6051808>0$ for all $x\geq 0$ and hence, $p(x)>0$ for all $x\geq 50$ as desired. [^19] One can check that p(49)=303155<0p(49)=-303155<0. This is why we addressed $4\leq m\leq 49$ and $m\geq 50$ separately.
Next, we prove that $r_{n},r^{\prime}_{n}$ are the only two positive roots for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us recall Descartes’ rule of signs. It says that for a polynomial $a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{d}x^{d}$ of degree $d$, the number $N$ of positive roots and the number $N^{\prime}$ of sign changes in the coefficients have the same parity and $N\leq N^{\prime}$. The number $N^{\prime}$ is the number of indices $0\leq j\leq d-1$ such that $a_{j}a_{j+1}<0$. The rest of proof is devoted to showing that there are only two sign changes in the coefficients of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for every even integer $n\geq 8$. Since we already showed the existence of at least two positive roots $r_{n},r^{\prime}_{n}$, it follows by Descartes’ rule of signs that $r_{n},r^{\prime}_{n}$ are the only two positive roots for $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$. Let us first determine the coefficients of $P_{n}^{(1)}(r)$ and $P_{n}^{(2)}(r)$. In case of $P_{n}^{(1)}(r)$, they are given by
$P_{n}^{(1)}(r)=\sum_{j=0}^{\frac{n}{2}}a^{(1)}_{j}r^{j},\,\,\,\,\,a^{(1)}_{j}=\frac{{\frac{n}{2}\choose j}}{n^{j}}.$ (189)
In case of $P_{n}^{(2)}(r)$, we write
$P_{n}^{(2)}(r)$ $=$ $\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-1}\frac{1}{2^{i}i!}\sum_{j=0}^{i}{i\choose j}r^{j}$ (190)
$=$ $\frac{1}{\sqrt{e}}\sum_{j=0}^{\frac{n}{2}-1}\sum_{i=j}^{\frac{n}{2}-1}\frac{{i\choose j}}{2^{i}i!}r^{j},$
where in the second step we have changed the order of summations. This shows that
$P_{n}^{(2)}(r)=\sum_{j=0}^{\frac{n}{2}}a^{(2)}_{j}r^{j},\,\,\,\,\,a^{(2)}_{j}=\left\{\begin{array}[]{cc}\frac{1}{\sqrt{e}}\sum_{i=j}^{\frac{n}{2}-1}\frac{{i\choose j}}{2^{i}i!}&0\leq j\leq\frac{n}{2}-1\\ 0&j=\frac{n}{2}\end{array}\right.$. (193)
The coefficients of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ are $a_{j}^{(1)}-a_{j}^{(2)}$ for $j=0,1,2,\cdots,\frac{n}{2}$. These coefficients incur a total of two sign changes. More precisely, we show
$a^{(1)}_{0}-a^{(2)}_{0}>0,\,\,\,\,\,a^{(1)}_{1}-a^{(2)}_{1}>0,$ (194)
$a^{(1)}_{j}-a^{(2)}_{j}<0,\,\,\,\,\,j=2,3,\cdots,\frac{n}{2}-1$ (195)
and
$a^{(1)}_{\frac{n}{2}}-a^{(2)}_{\frac{n}{2}}>0.$ (196)
Verifying (194) is trivial. In fact, $a_{0}^{(1)}=P_{n}^{(1)}(0),a_{0}^{(2)}=P_{n}^{(2)}(0)$ and we checked in (174) that $P_{n}^{(1)}(0)>P_{n}^{(2)}(0)$. Also,
$a^{(1)}_{1}=\frac{\frac{n}{2}}{n}=\frac{1}{2}$ (197)
and
$a^{(2)}_{1}$ $=$ $\frac{1}{\sqrt{e}}\sum_{i=1}^{\frac{n}{2}-1}\frac{i}{2^{i}i!}=\frac{1}{\sqrt{e}}\sum_{i=1}^{\frac{n}{2}-1}\frac{1}{2^{i}(i-1)!}$ (198)
$=$ $\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-2}\frac{1}{2^{i+1}i!}=\frac{1}{2\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-2}\frac{1}{2^{i}i!}<\frac{1}{2\sqrt{e}}\sum_{i=0}^{\infty}\frac{(\frac{1}{2})^{i}}{i!}=\frac{1}{2}.$
The inequality in (196) is also trivial as $a_{\frac{n}{2}}^{(2)}=0<a_{\frac{n}{2}}^{(1)}=\frac{1}{n^{\frac{n}{2}}}$. To prove (195), first we write it in a different form. The condition $a^{(1)}_{j}<a^{(2)}_{j}$ for $j=2,3,\cdots,\frac{n}{2}-1$ can be written as
$\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=j}^{\frac{n}{2}-1}\frac{{i\choose j}}{2^{i}i!}$ (199)
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ $\frac{{\frac{n}{2}\choose j}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{{i+j\choose j}}{2^{i+j}(i+j)!}$
$\iff$ $\frac{\frac{\frac{n}{2}!}{j!(\frac{n}{2}-j)!}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{\frac{(i+j)!}{j!i!}}{2^{i+j}(i+j)!}$
$\iff$ $\frac{\frac{\frac{n}{2}!}{(\frac{n}{2}-j)!}}{n^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{1}{2^{i+j}i!}$
$\iff$ $\frac{\frac{\frac{n}{2}!}{(\frac{n}{2}-j)!}}{(\frac{n}{2})^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{1}{2^{i}i!}$
$\iff$ $\frac{\frac{n}{2}(\frac{n}{2}-1)(\frac{n}{2}-2)\cdots(\frac{n}{2}-j+1)}{(\frac{n}{2})^{j}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{\frac{n}{2}-j-1}\frac{1}{2^{i}i!},$
where in ()(*) we have changed the index $i$ to $i+j$. Call $m=\frac{n}{2}$. Then we need to show
$\frac{(m-1)(m-2)\cdots(m-j+1)}{m^{j-1}}<\frac{1}{\sqrt{e}}\sum_{i=0}^{m-j-1}\frac{1}{2^{i}i!},\,\,\,\,\,2\leq j\leq m-1,$ (200)
holds for all $m\geq 4$. We address $m=4$ and $m\geq 5$ separately.
  1. 1.
    Let $m=4$. The possibilities for $j$ are $j=2,3$. If $j=2$, then (200) becomes $\frac{3}{4}<\frac{1}{\sqrt{e}}(1+\frac{1}{2})$ which is true. If $j=3$, then (200) becomes $\frac{3\times 2}{16}<\frac{1}{\sqrt{e}}\times 1$ which is also true.
  2. 2.
    Assume $m\geq 5$. Let $\boldsymbol{x}\sim\mathrm{Poisson}(\frac{1}{2})$ be a Poisson random variable with parameter $\frac{1}{2}$. Then the right side of (200) can be written as
    $\frac{1}{\sqrt{e}}\sum_{i=0}^{m-j-1}\frac{1}{2^{i}i!}$ $=$ $\sum_{i=0}^{m-j-1}\frac{e^{-\frac{1}{2}}(\frac{1}{2})^{i}}{i!}$ (201)
    $=$ $\Pr(\boldsymbol{x}\leq m-j-1)$
    $=$ $1-\Pr(\boldsymbol{x}\geq m-j)$
    $\stackrel{{\scriptstyle(a)}}{{\geq}}$ $1-\frac{\mathbb{E}[\boldsymbol{x}]}{m-j}$
    $\stackrel{{\scriptstyle(b)}}{{=}}$ $1-\frac{\frac{1}{2}}{m-j}$
    $=$ $\frac{m-j-\frac{1}{2}}{m-j},$
    where $(a)$ is due to Markov’s inequality and $(b)$ is due to $\mathbb{E}[\boldsymbol{x}]=\frac{1}{2}$. By (200) and (201), it is enough to show that for every $m\geq 5$,
    $\frac{(m-1)(m-2)\cdots(m-j+1)}{m^{j-1}}<\frac{m-j-\frac{1}{2}}{m-j},\,\,\,\,\,j=2,3,\cdots,m-1.$ (202)
    For $j=2$, (202) becomes $\frac{m-1}{m}<\frac{m-\frac{5}{2}}{m-2}$ which is equivalent to $(m-1)(m-2)<m(m-\frac{5}{2})$. This simplifies to $m>4$ which is true. [^20] This is why we separated the cases $m=4$ and $m>4$. If $j\geq 3$, we write (202) as
    $\frac{(m-1)(m-2)\cdots(m-j+2)}{m^{j-2}}\times\frac{m-j+1}{m}<\frac{m-j-\frac{1}{2}}{m-j},$ (203)
    or equivalently,
    $\Big(1-\frac{1}{m}\Big)\Big(1-\frac{2}{m}\Big)\cdots\Big(1-\frac{j-2}{m}\Big)\frac{m-j+1}{m}<\frac{m-j-\frac{1}{2}}{m-j}.$ (204)
    Since $(1-\frac{1}{m})(1-\frac{2}{m})\cdots(1-\frac{j-2}{m})<1$, it suffices to prove
    $\frac{m-j+1}{m}<\frac{m-j-\frac{1}{2}}{m-j},\,\,\,\,\,j=3,\cdots,m-1.$ (205)
    This is equivalent to $(m-j+1)(m-j)<m(m-j-\frac{1}{2})$ which simplifies to
    $m>\frac{j(j-1)}{j-\frac{3}{2}},\,\,\,\,\,j=3,\cdots,m-1.$ (206)
    The function $f(x)=\frac{x(x-1)}{x-\frac{3}{2}}$ is monotone increasing for $x\geq 3$. Therefore, (206) holds if and only if $m>f(m-1)=\frac{(m-1)(m-1-1)}{m-\frac{3}{2}}$. This simplifies to $m>\frac{4}{3}$ which is true. The proof of Proposition 5 is now complete.
X Concluding Remarks
We made progress in improving HWI by increasing the absolute constant $\kappa$ in its formula from the best currently known value of $0.125$ to at least $0.145$ in the general symmetric case. This was achieved by determining the best quadratic upper bound on the function $\ln(1-x)$ around $x=0$. In the positive-semidefinite case, it was shown that LMI implies HWI with $\kappa=1-\frac{\sqrt{3}}{2}$. We verified that indeed a larger value $\kappa=\frac{9-\sqrt{17}}{32}$ is possible through improving LMI. This improvement came about in two forms called the augmented LMI and the optimal LMI. The latter requires computing the unique root of a quintic polynomial in the interval $(\frac{2}{3},1)$ which must be done numerically. The former approximates this unique root by a closed-form expression where the underlying approximation error was demonstrated to be less than $0.035$ regardless of the positive-semidefinite matrix $A$ and the tail parameter $t$.
We explored beyond HWI for general symmetric matrix $A$ by identifying the best polynomial upper bound of arbitrary degree $m+1$ on the function $\ln(1-x)$ around $x=0$ where $m\geq 1$. As a result, a sequence of concentration bounds indexed by $m$ was developed where the case $m=1$ recovers HWI. These bounds are written in terms of Schatten norms of $A$ and it was conjectured that they undergo a phase transition in the sense that for given $A$, there exists a threshold $\tau_{c}$ such that if $t<\tau_{c}$, then $m=1$ (HWI) is the tightest and if $t>\tau_{c}$, then letting $m$ grow to infinity results in the sharpest bound. This led to what we called the $m_{\infty}$-bound which admits a very simple formula and it depends on $A$ only through the dimension $n$ and its operator norm $\|A\|$. An estimate (upper bound) was also derived on $\tau_{c}$.
As mentioned earlier, every upper bound of index $m\geq 2$ in that derived sequence is looser than HWI if $t$ is sufficiently small. We raised the question whether one can improve upon HWI for all sufficiently small tail parameter $t$. This question was answered positively by finding a concentration bound that depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$. It was called a twin to HWI due to the fact that its exponent scales similarly to the exponent in HWI, however, with a larger coefficient such that it outperforms HWI for both small and large values of $t$.
Finally, we considered a scenario where the positive-semidefinite matrix $A$ is known only through the dimension $n$ and its largest eigenvalue (operator norm). We presented five upper bounds on the tail probability of the Gaussian quadratic chaos, namely, the $m_{\infty}$-bound, relaxed versions of HW and LM bounds, the $\chi^{2}$-bound and the large deviations bound. All of these bounds depend on $A$ and $t$ through the ratio $\frac{t}{\|A\|}$. It was proved that the $m_{\infty}$-bound is sharper than the relaxed HW and LM bounds for every $n\geq 1$, $A$ and $t>0$. The $m_{\infty}$-bound is also sharper than the large deviations bound for every $n\geq 2$, $A$ and $t>0$. If $n=2,4,6$, then the $\chi^{2}$-bound is tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, there exist positive numbers $r_{n},r^{\prime}_{n}$ such that the $m_{\infty}$-bound is tighter than the $\chi^{2}$-bound if and only if $r_{n}<\frac{t}{\|A\|}<r^{\prime}_{n}$. Moreover, as the dimension $n$ grows, $r_{n}$ moves towards zero and $r^{\prime}_{n}$ grows to infinity. It was also shown that the so-called large deviations upper bound is too loose to be useful in the sense that it is weaker than the $m_{\infty}$-bound for every $n\geq 2$, $A$ and $t>0$.
Appendix A; Proof of Lemma 1
Define the function $f$ by
$f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+a|x|^{m+1}+\ln(1-x).$ (207)
We study the cases $x>0$ and $x<0$ separately.
  1. 1.
    Let $x>0$. Then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ and we get
    $f^{\prime}(x)$ $=$ $1+x+x^{2}+\cdots+x^{m-1}+(m+1)ax^{m}-\frac{1}{1-x}$ (208)
    $=$ $\frac{1-x^{m}+(m+1)ax^{m}(1-x)-1}{1-x}$
    $=$ $\frac{x^{m}((m+1)a-1-(m+1)ax)}{1-x}.$
    The only first-order critical numbers for $f$ are 0 and $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, then one easily checks that $f^{\prime}(x)<0$ over the open interval $(0,1)$. Since $f(0)=0$, the Mean Value Theorem (MVT) implies that $f(x)<0$ over $(0,1)$ and hence, (10) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $0<x_{0}<1$, $f^{\prime}(x)>0$ over the interval $(0,x_{0})$ and $f^{\prime}(x)<0$ over the interval $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=$-$\infty$. This tells us that $f$ rises above 0 on right of $x=0$, reaches a maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to -$\infty$ on left of $x=1$. As such, $f$ has a unique zero (x-crossing) somewhere over the interval $(x_{0},1)$. We let $b$ be this zero. Then (23) clearly holds for $0\leq x\leq b$. Moreover, the equation $f(b)=0$ gives $a=\theta_{m}(b)$ as defined in (11).
  2. 2.
    Let $x<0$. If $m$ is odd, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ as in the previous case and one easily checks that for $a>\frac{1}{m+1}$, $f^{\prime}(x)$ given in (208) is negative for every $x<0$. Then MVT rules that $f(x)>0$ for every $x<0$ as desired. If $m$ is even, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}-ax^{m+1}+\ln(1-x)$. A similar computation as in (208) gives that
    $f^{\prime}(x)=\frac{-x^{m}((m+1)a+1-(m+1)ax)}{1-x}.$ (209)
    Since $m$ is even, we have $f^{\prime}(x)<0$ for every $x<0$ and MVT applies once again to confirm that $f(x)>0$ for $x<0$.
Appendix B; Proof of Lemma 2
Assume $m\geq 1$ is an integer and fix $\frac{m+1}{m+2}<a<1$. Define the function $f$ by
$\hskip 28.45274ptf(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+\frac{x^{m+1}}{(m+1)(1-ax)}+\ln(1-x),\,\,\,\,\,0\leq x<1.$ (210)
One can easily check that
$f^{\prime}(x)=\frac{x^{m+1}\big((m+2)a-m-1-a((m+1)a-m)x\big)}{(m+1)(1-x)(1-ax)^{2}}.$ (211)
Thanks to the condition $\frac{m+1}{m+2}<a<1$, the only positive first-order critical number for $f$ over the interval $(0,1)$ is given by $x_{0}=\frac{m+1-(m+2)a}{a(m-(m+1)a)}$. Moreover, $f^{\prime}(x)>0$ for $0<x<x_{0}$ and $f^{\prime}(x)<0$ for $x_{0}<x<1$. Since $f(0)=0$, then MVT implies that $f(x)>0$ for every $x$ in the interval [^21] In fact, $f$ has a unique x-crossing at some number $x_{1}$ inside the interval $(x_{0},1)$ and $f(x)\geq 0$ for every $0\leq x\leq x_{1}$. Hence, (119) also holds with $b=x_{1}$. However, unlike $x_{0}$, the number $x_{1}$ does not admit a closed-form expression. $(0,x_{0}]$ and hence, (119) holds with $b=x_{0}$.
Appendix C; The choice of $a=a_{0}$ in (128) yields the largest $\kappa$
In order to compare $\Lambda(t,a)$ in (20) with $\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}$, we address the cases $c\leq 1$ and $c>1$ separately.
  1. 1.
    Let $c\leq 1$, or equivalently, $\frac{2}{3}<a\leq a_{0}$. Then
    $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big(\sqrt{1+2a\rho}-1\big)&0<\rho\leq c\\ \frac{b}{2\rho}-\frac{b^{2}}{4(1-ab)\rho^{2}}&c<\rho\leq 1\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>1\end{array}\right.$. (215)
    One can easily check that this is a continuous function of $\rho$ and it achieves its absolute minimum value of $\frac{b}{2}-\frac{b^{2}}{4(1-ab)}=\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}$ at $\rho=1$.
  2. 2.
    Let $c>1$, or equivalently, $a_{0}<a<1$. Then
    $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big(\sqrt{1+2a\rho}-1\big)&0<\rho\leq 1\\ \frac{1}{2a}-\frac{1}{2a^{2}\rho}(\sqrt{1+2a\rho}-1)&1<\rho\leq c\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>c\end{array}\right.$. (219)
    Once again, this is a continuous function of $\rho$ and it achieves its absolute minimum value of $\frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)$ at $\rho=1$.
We have shown that
$\min_{\rho>0}-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}&\frac{2}{3}<a\leq a_{0}\\ \frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)&a_{0}<a<1\end{array}\right.$. (222)
We need to choose $\frac{2}{3}<a<1$ such that the right hand side in (222) is as large as possible. Fig. 12 shows the graph of this function of $a$ over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
Figure 12: The graph of the function of $a$ that sits on the right side of (222) over the interval $(\frac{2}{3},1)$. It achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
Appendix D; Proof of Lemma 3
Tedious algebra shows that
$-ux+vx^{2}+wx^{3}=\frac{g(u,v,w)}{27w^{\frac{1}{2}}},$ (223)
where $g(u,v,w)$ is defined by
$g(u,v,w)=\frac{1}{w^{\frac{3}{2}}}(\sqrt{v^{2}+3uw}-v)(v\sqrt{v^{2}+3uw}-v^{2}-6uw).$ (224)
We show that $g$ is decreasing in $w$ for given $u,v$. Another round of tedious algebra shows
$\frac{\partial g}{\partial w}=-\frac{3v}{2w^{\frac{5}{2}}}\big(2v^{2}+3uw-2v\sqrt{v^{2}+3uw}\big).$ (225)
Then $\frac{\partial g}{\partial w}\leq 0$ follows by noting that $2v^{2}+3uw\geq 2v\sqrt{v^{2}+3uw}$ for all $u,v,w>0$.
Appendix E; Proof of (36)
For an even integer $n\geq 8$, let us denote $m=\frac{n}{2}\geq 4$. During the proof, we adopt the notation $\forall_{\infty}m$ to say that a statement indexed by the integer $m$ holds for all but finitely many values for $m$, i.e., it eventually holds for all sufficiently large $m$. We begin by showing that
$\forall_{\infty}m\geq 4,\,\,\,r^{\prime}_{2m}>2m.$ (226)
This establishes the second claim in (36). Recall from the proof of Proposition 5 that $r_{n},r^{\prime}_{n}$ are the only two positive zeros for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ and that this polynomial is negative if and only if $r_{n}<r<r^{\prime}_{n}$. In order to verify (226), it is enough to show that
$\forall_{\infty}m\geq 4,\,\,\,P_{2m}^{(1)}(2m)<P_{2m}^{(2)}(2m).$ (227)
We have $P_{2m}^{(2)}(2m)=\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{2^{i}i!}(1+2m)^{i}$. We show that the last term in this sum is eventually larger than $P_{2m}^{(1)}(2m)=2^{m}$, i.e.,
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{(1+2m)^{m-1}}{2^{m-1}(m-1)!}=\frac{1}{\sqrt{e}}\frac{(m+\frac{1}{2})^{m-1}}{(m-1)!}.$ (228)
This is easily done by invoking the Stirling’s approximation formula. Let us recall the inequality [^22] See Section 2.9 in [11].
$k!<\sqrt{2\pi}\,k^{k+\frac{1}{2}}e^{-k+\frac{1}{12k}},\,\,\,k\geq 1,$ (229)
due to Feller [11, 12]. Plugging $k=m-1\geq 3$ in (229) and observing that $\sqrt{2\pi}e^{\frac{1}{12(m-1)}}\leq\sqrt{2\pi}e^{\frac{1}{36}}\approx 2.577<e$, we get $(m-1)!<(m-1)^{m-\frac{1}{2}}e^{2-m}$ and hence,
$\frac{e^{m-2}}{(m-1)^{m-\frac{1}{2}}}<\frac{1}{(m-1)!},\,\,\,m\geq 4.$ (230)
By (230), the proof of (228) is complete if we can verify the stronger statement
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{(m+\frac{1}{2})^{m-1}e^{m-2}}{(m-1)^{m-\frac{1}{2}}}.$ (231)
Note that $\frac{(m+\frac{1}{2})^{m-1}}{(m-1)^{m-\frac{1}{2}}}>\frac{(m+\frac{1}{2})^{m-1}}{(m+\frac{1}{2})^{m-\frac{1}{2}}}=\frac{1}{(m+\frac{1}{2})^{\frac{1}{2}}}$. Thus, (231) follows if
$\forall_{\infty}m\geq 4,\,\,\,2^{m}<\frac{1}{\sqrt{e}}\frac{e^{m-2}}{(m+\frac{1}{2})^{\frac{1}{2}}}.$ (232)
But, this is certainly true as the right hand side scales like $e^{m-O(\frac{1}{2}\ln m)}$. The proof of (226) is now complete.
To verify the first claim in (36), we show that
$\forall\epsilon>0,\,\,\,\forall_{\infty}m\geq 4,\,\,\,r_{2m}<\epsilon.$ (233)
Following the discussion before (227), it suffices to prove that
$\forall_{\infty}m\geq 4,\,\,\,P_{2m}^{(1)}(\epsilon)<P_{2m}^{(2)}(\epsilon).$ (234)
The proof follows similar lines of reasoning to show $P_{n}^{(1)}(1)<P_{n}^{(2)}(1)$ in Lemma 4. Fix $\epsilon>0$. Let $\boldsymbol{x}\sim\mathrm{Poisson}(\frac{1+\epsilon}{2})$ be a Poisson random variable with parameter $\frac{1+\epsilon}{2}$. Then
$P_{2m}^{(2)}(\epsilon)$ $=$ $\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{(1+\epsilon)^{i}}{2^{i}i!}$ (235)
$=$ $e^{\frac{\epsilon}{2}}\sum_{i=0}^{m-1}\frac{e^{-\frac{1+\epsilon}{2}}(\frac{1+\epsilon}{2})^{i}}{i!}$
$=$ $e^{\frac{\epsilon}{2}}\Pr(\boldsymbol{x}\leq m-1)$
$=$ $e^{\frac{\epsilon}{2}}(1-\Pr(\boldsymbol{x}\geq m))$
$\stackrel{{\scriptstyle(a)}}{{\geq}}$ $e^{\frac{\epsilon}{2}}\Big(1-\frac{\mathbb{E}[\boldsymbol{x}^{2}]}{m^{2}}\Big)$
$\stackrel{{\scriptstyle(b)}}{{=}}$ $e^{\frac{\epsilon}{2}}\Big(1-\frac{\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}}{m^{2}}\Big)$
$=$ $e^{\frac{\epsilon}{2}}\Big(1-\frac{\epsilon^{2}+4\epsilon+3}{4m^{2}}\Big),$
where $(a)$ follows by Markov’s inequality and $(b)$ is due to $\mathbb{E}[\boldsymbol{x}^{2}]=\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}$. Also, $P_{2m}^{(1)}(\epsilon)=(1+\frac{\epsilon}{2m})^{m}$. By (235), if we can show that
$\forall_{\infty}m\geq 4,\,\,\,\Big(1+\frac{\epsilon}{2m}\Big)^{m}<e^{\frac{\epsilon}{2}}\Big(1-\frac{\epsilon^{2}+4\epsilon+3}{4m^{2}}\Big),$ (236)
then (234) follows. Define the function
$f_{\epsilon}(x)=\frac{\epsilon}{2}+\ln\Big(1-\frac{\epsilon^{2}+4\epsilon+3}{4x^{2}}\Big)-x\ln\Big(1+\frac{\epsilon}{2x}\Big),\,\,\,x>\frac{1}{2}\sqrt{\epsilon^{2}+4\epsilon+3}.$ (237)
The statement in (236) can be written as $\forall_{\infty}m\geq 4,e^{f_{\epsilon}(m)}>1$. Thus, it is enough to prove that $f_{\epsilon}(x)>0$ for sufficiently large $x$. This is accomplished if we can verify
$\lim_{x\to\infty}f_{\epsilon}(x)=0$ (238)
and
$f^{\prime}_{\epsilon}(x)<0,$ (239)
for all sufficiently large $x$. Checking (238) is trivial. To verify (239), we use a similar argument and show that
$\lim_{x\to\infty}f^{\prime}_{\epsilon}(x)=0$ (240)
and
$f^{\prime\prime}_{\epsilon}(x)>0,$ (241)
for all sufficiently large $x$. We have
$f^{\prime}_{\epsilon}(x)=\frac{2(\epsilon^{2}+4\epsilon+3)}{x(4x^{2}-\epsilon^{2}-4\epsilon-3)}+\frac{\epsilon}{2x+\epsilon}-\ln\Big(1+\frac{\epsilon}{2x}\Big).$ (242)
Then (240) clearly holds. Also,
$f^{\prime\prime}_{\epsilon}(x)=\frac{p_{\epsilon}(x)}{x^{2}(2x+\epsilon)^{2}(4x^{2}-\epsilon^{2}-4\epsilon-3)^{2}},$ (243)
where $p_{\epsilon}(x)$ is a polynomial of degree $5$ in $x$ given by [^23] The coefficients behind powers of $x$ in $p_{\epsilon}(x)$ are all polynomials in $\epsilon$.
$p_{\epsilon}(x)=16\epsilon^{2}x^{5}+\textrm{terms of lower degrees in $x$}.$ (244)
The leading term in the polynomial $p_{\epsilon}(x)$ is $16\epsilon^{2}x^{5}$. Since the coefficient $16\epsilon^{2}$ behind $x^{5}$ is positive, then (241) must hold for all sufficiently large $x$ depending on $\epsilon$ of course. The proof of (233) is complete.